### v1

# Think in Safety: Unveiling and Mitigating Safety Alignment Collapse in Multimodal Large Reasoning Model

Abstract
The rapid development of multimodal large reasoning models (MLRMs) has demonstrated broad application potential, yet their safety and reliability remain critical concerns that require systematic exploration. To address this gap, we conduct a comprehensive and systematic safety evaluation of 11 MLRMs across 5 benchmarks and unveil prevalent safety degradation phenomena in most advanced models. Moreover, our analysis reveals distinct safety patterns across different benchmarks: significant safety degradation is observed across jailbreak robustness benchmarks, whereas safety-awareness benchmarks demonstrate less pronounced degradation. In particular, a long thought process in some scenarios even enhances safety performance. Therefore, it is a potential approach to addressing safety issues in MLRMs by leveraging the intrinsic reasoning capabilities of the model to detect unsafe intent. To operationalize this insight, we construct a multimodal tuning dataset that incorporates a safety-oriented thought process. Experimental results from fine-tuning existing MLRMs with this dataset effectively enhances the safety on both jailbreak robustness and safety-awareness benchmarks. This study provides a new perspective for developing safe MLRMs. Our dataset is available at https://github.com/xinyuelou/Think-in-Safety.
Warning: this paper contains example data that may be offensive or harmful.
Think in Safety: Unveiling and Mitigating Safety Alignment Collapse in Multimodal Large Reasoning Model
Xinyue Lou1111Equal contribution. Contact e-mail: louxinyue@bjtu.edu .cn, youlee@bjtu.edu.cn, You Li1111Equal contribution. Contact e-mail: louxinyue@bjtu.edu .cn, youlee@bjtu.edu.cn, Jinan Xu1, Xiangyu Shi1, Chi Chen2, Kaiyu Huang1 1 Beijing Jiaotong University, Beijing, China 2 Tsinghua University, Beijing, China
1 Introduction
The development of large reasoning models (LRMs) such as OpenAI o1 Jaech et al. (2024) and DeepSeek-R1 Guo et al. (2025) has unveiled significant potential in addressing complex reasoning tasks Chan et al. (2024); Chen et al. (2024), including mathematical problem-solving Shao et al. (2024) and code generation El-Kishky et al. (2025). Consequently, the reasoning paradigm has been naturally extended to multimodal scenarios, giving rise to multimodal large reasoning models (MLRMs) Peng et al. (2025a); Yao et al. (2024); Team et al. (2025). The MLRMs are capable of generating long reasoning pathways and exhibit remarkable spatial reasoning abilities, thereby broadening their real-world applicability across diverse scenarios Zhang et al. (2025a); Huang et al. (2025); Wu et al. (2025).
Figure 1: Examples of multimodal safety benchmarks and their corresponding responses on different models.
While LRMs have led to a transformative leap in reasoning performance, prior studies have indicated that the exposure of the chain-of-thought (CoT) process inadvertently undermines safety, as adversaries exploit intermediate reasoning steps to circumvent alignment safeguards Jiang et al. (2025); Zhou et al. (2025); Ying et al. (2025). This vulnerability has also been observed in MLRMs Fang et al. (2025). However, current research on the safety of MLRMs remains limited, primarily focusing on two jailbreak robustness benchmarks, with insufficient exploration of the broader spectrum of safety challenges in multimodal settings.
As shown in Figure 1, we categorize the safety benchmarks of multimodal models into two key aspects: safety-awareness and jailbreak robustness. Jailbreak robustness benchmarks focus on evaluating the resilience of the model against deliberately crafted or modified textual prompts and visual inputs that aim to bypass established safety defense mechanisms. Safety-awareness benchmarks emphasize the capability of models to proactively identify potential safety risks embedded in user inputs, aligning more closely with the complex and dynamic safety demands encountered in real-world applications. Compared to unimodal LRMs, MLRMs exhibits the following research questions:
RQ1: How does MLRMs affect safety on different types of benchmarks compared with Multimodal Large Language Models (MLLMs)?
RQ2: What are the limitations of incorporating additional modalities in MLRMs for safety concerns?
RQ3: What is the impact of reasoning pathways in MLRMs on the safety performance?
To investigate the above research questions and challenges, in this study, we first conduct a systematic safety evaluation of advanced MLRMs, including Kimi-VL-thinking Team et al. (2025), R1-Onevision Yang et al. (2025) and etc. The results demonstrate that MLRMs have significant negative impact on the safety performance, while the performance degradation is task-dependent. In particular, MLRMs could identify more potential safety risks through deliberate thinking, leading to higher safety scores on safety-awareness benchmarks, which provides a novel perspective for mitigating safety degradation in MLRMs. Motivated by this, we further propose a data construction method that incorporates safety-oriented thought process to investigate the effectiveness of this insight.
To sum up, our contributions are summarized as follows:
  • This study conducts a systematic evaluation of MLRMs and investigates the empirical results, revealing several novel findings and providing new perspectives for the development of safer MLRMs.
  • We construct a multimodal fine-tuning dataset with safety-oriented thought process for safety alignment, alleviating the issue of incorporating additional modalities.
  • Experimental results demonstrate that our method improve the safety performance of MLRMs across multiple benchmarks by enabling self-correction thinking along the reasoning pathways, compared with previous defense methods.
Model Name Base Model
LlamaV-o1 Thawakar et al. (2025) Llama3.2-vision-11B
Llava-cot Xu et al. (2024) Llama3.2-vision-11B
Mulberry-Llama Yao et al. (2024) Llama3.2-vision-11B
LMM-R1 Peng et al. (2025b) Qwen2.5-VL-3B
R1-Onevision Yang et al. (2025) Qwen2.5-VL-7B
MM-Eureka Meng et al. (2025) InternVL-2.5-8B
Kimi-VL-thinking Team et al. (2025) Kimi-VL-instruct
Skywork-R1V Peng et al. (2025a) -
QVQ-Preview -
Gemini2.0-flash-thinking -
Claude3.7-sonnet-thinking -
Table 1: List of the evaluated MLRMs and their corresponding base models.
Jailbreak Robustness Safety-Awareness
[…]
Gemini2.0-flash-thinking 89.80 73.48 50.36 39.16 57.32
Claude3.7-sonnet-thinking 32.20 33.94 0.49 59.24 57.76
QVQ-Preview 70.80 69.29 37.86 34.13 50.68
Skywork-R1V 85.80 72.68 35.28 35.33 50.42
Llama3.2-vision-11B(base) 55.80 38.45 5.71 37.13 52.26
LlamaV-o1 59.40(+3.60) 53.93(+15.48) 13.57(+7.86) 33.93(-3.20) 51.59(-0.67)
Llama3.2-vision-11B(base) 55.80 38.45 5.71 37.13 52.26
Llava-cot 84.80(+29.00) 72.26(+33.81) 33.57(+27.86) 26.95(-10.18) 51.26(-1.00)
Llama3.2-vision-11B(base) 55.80 38.45 5.71 37.13 52.26
Mulberry-Llama 67.40(+11.60) 64.70(+26.25) 13.21(+7.50) 37.72(+0.59) 54.09(+1.83)
Qwen2.5-VL-3B(base) 66.27 66.18 12.14 24.55 52.35
LMM-R1 69.80(+3.53) 68.15(+1.97) 18.21(+6.07) 21.56(-2.99) 53.02(+0.67)
Qwen2.5-VL-7B(base) 67.20 66.49 12.50 29.94 51.09
R1-Onevision 72.20(+5.00) 79.57(+13.08) 32.14(+19.64) 17.31(-12.63) 49.69(-1.40)
InternVL-2.5-8B(base) 71.40 59.64 15.00 28.14 50.84
$M$-Eureka 72.20(+0.80) 60.12(+0.48) 11.79(-3.21) 28.14(-0.00) 50.59(-0.25)
Kimi-VL-instruct(base) 80.40 47.74 22.50 25.00 50.44
Kimi-VL-thinking 87.00(+6.60) 61.49(+13.75) 33.93(+11.43) 35.93(+10.93) 51.42(+0.98)
Table 2: Variation of safety performance for MLRMs across various benchmarks. $\downarrow$ means the lower score the safer, while $\uparrow$ means the higher the better. The safety performance variation has been marked in brackets, where the red color represents safety deterioration and green stands for safety improvement.
2 Safety Evaluation of MLRMs
2.1 Evaluation Settings
Datasets.
To comprehensively assess the safety performance of MLRMs across diverse scenarios, we adopt benchmark datasets from two distinct perspectives: safety-awareness and jailbreak robustness. For the assessment of safety-awareness, we employ SIUO Wang et al. (2025) and MSSBench Zhou et al. (2024) datasets, while for evaluating jailbreak robustness, we employ MMSafetyBench Liu et al. (2024b), FigStep Gong et al. (2025) and JailBreaKV Luo et al. (2024).
In safety-awareness tasks, models need to reason jointly over both visual and textual inputs to infer user intent, identify potential safety risks, and assess whether the input should be treated as safe or unsafe. These tasks pose significant challenges to the multimodal reasoning and safety alignment capabilities of models. In contrast, jailbreak robustness benchmarks involve adversarial attacks, such as the inclusion of malicious crafted prompt, aimed at circumventing the safety constraints of models. Given that these two task categories examine safety alignment from fundamentally different perspectives, we analyze the results separately. Further details regarding the datasets are provided in Appendix A.
Models and Configurations.
We evaluate a total of 11 MLRMs, including both proprietary models and open-source models along with their corresponding base models, such as Kimi-VL-thinking Team et al. (2025), R1-Onevision Yang et al. (2025), Mulberry Yao et al. (2024), and LlamaV-o1 Thawakar et al. (2025). Detailed information of the models are provided in Table 1. For reasoning models that have undergone extensive fine-tuning—via either supervised learning or reinforcement learning—we additionally evaluate their corresponding base models. This approach allows us to trace safety-related changes across model variants. All experiments were conducted using four NVIDIA A100-80G GPUs.
Figure 2: Case study of the better safety consideration on safety-awareness tasks. Kimi-VL-Instruct directly outputs the answer that ignores the potential risk, while Kimi-VL-Thinking dives deeper into the insidious safety issue with stronger reasoning abilities.
Metrics and Evaluator.
For jailbreak robustness benchmarks, we adopt the standard Attack Success Rate (ASR) metric, which quantifies the percentage of instances where the model produces harmful outputs with its safety mechanisms circumvented. A lower ASR indicates stronger safety alignment. For safety-awareness benchmarks, we follow established evaluation protocols to compute a safety score, where higher scores denote better safety performance. Further evaluation settings are detailed in Appendix A. Following the settings of MMSafetyBench, MSS, and SIUO utilizing OpenAI API, we employ GPT-4o-mini Hurst et al. (2024) as the evaluator and adopt identical configurations to ensure consistency across assessments.
2.2 Safety Results and Inspection
Finding 1 ($\Rightarrow$ RQ1): The safety performance of MLRMs requires improvement and exhibits a notable security degradation phenomenon.
Overall Performance.
As shown in Table 2, existing multimodal reasoning models exhibit significant security vulnerabilities, with most open-source models performing unsatisfactorily across various benchmarks. In particular, Kimi-VL-thinking, Skywork-R1V and Llava-cot successfully defend against only about 15% of malicious queries on Figstep. This finding highlights the necessity of ensuring safety alignment alongside improvements in multimodal reasoning capabilities.
Security Degradation.
Most reasoning models exhibit a notable increase in ASR on jailbreak robustness benchmarks relative to their corresponding foundation models, with an average increase of approximately 11.53%. This suggests that the internal safety alignment mechanisms are compromised during the process of enhancing reasoning capabilities, thereby rendering the models more susceptible to jailbreak attacks. In contrast, the degradation of safety performance is less evident in safety-awareness benchmarks. In some cases, models even demonstrate improved safety performance after engaging in extended thought process. A detailed analysis of this phenomenon is presented in Finding 2.
Finding 2 ($\Rightarrow$ RQ1): MLRM paradoxically enhances safety performance via long reasoning on safety-awareness benchmarks, due to improved capability in identifying unsafe intent.
Performance on Safety-Awareness.
As shown in Table 2, the safety score of models such as Kimi-VL-thinking Team et al. (2025), LMM-R1 Peng et al. (2025b), and MM-Eureka Meng et al. (2025) are higher than their corresponding base models on safety-awareness benchmark. The effect is particularly pronounced for Kimi-VL-Thinking model, which achieves a 10.9% improvement on the SIUO benchmark. As illustrated in Figure 2, there is a difference in attention patterns according to the responses generated by these models and their corresponding base models. Reasoning models, when engaging in long reasoning, tend to focus more on potentially unsafe details present in both the visual and textual modalities. These models improve the safety performance through internal self-reflection and error-correction mechanisms. This encouraging finding offers a promising direction for mitigating safety degradation by leveraging the innate reasoning capabilities of the models themselves.
Figure 3: Examples of self-deception in responses generated by MLRMs. The content of reframing intent of users is highlighted with yellow background, and harmful content is marked with the red font.
[…]
Qwen2.5-VL-7B (I) 66.49 12.5 29.94
Qwen2.5-VL-7B (T) 48.21(-18.28) 6.79(-5.71) 38.10(+8.16)
R1-Onevision (I) 79.57 32.14 17.31
R1-Onevision (T) 65.02(-14.55) 23.21(-8.93) 23.35(+6.04)
Kimi-VL-instruct (I) 47.74 22.50 25.00
Kimi-VL-instruct (T) 57.56(+9.82) 19.29(-3.21) 25.00(+0.00)
Kimi-VL-thinking (I) 61.49 33.93 35.93
Kimi-VL-thinking (T) 55.71(-5.78) 29.64(-4.29) 29.34(-6.59)
Llama3.2-vision-11B (I) 38.45 5.71 37.13
Llama3.2-vision-11B (T) 58.52(+20.07) 44.29(+38.58) 22.50(-14.63)
Llava-cot (I) 72.26 33.57 26.95
Llava-cot (T) 60.06(-12.20) 40.36(+6.79) 28.14(+1.19)
InternVL2.5-8B (I) 59.64 15.00 28.14
InternVL2.5-8B (T) 48.57(-11.07) 9.64(-5.36) 27.55(-0.59)
MM-Eureka (I) 60.12 11.79 28.14
MM-Eureka (T) 50.24(-9.88) 10.71(-1.08) 31.14(+3.00)
Table 3: Safety performance when converting image into text caption. “I” indicates multimodal inputs, while “T” refers to text-only inputs.
Finding 3 ($\Rightarrow$ RQ2): The substitution of the image modality with text modality leads to a partial recovery of the safety defense capabilities of models.
Modality Ablation.
To investigate the impact of multimodal inputs on model safety, we convert the original multimodal inputs into text-only ones. In particular, we utilize the Qwen-2.5-VL-72B Bai et al. (2025) model to generate captions corresponding to the images, thereby replacing the visual information with its text representation. As shown in Table 3, most models demonstrate a noticeable improvement in safety performance when provided with unimodal (text-only) inputs, apart from the Llama-series models. This observation suggests that models are more effective at identifying harmful intent when reasoning over the text modality alone. In contrast, detecting harmful intent when reasoning over both image and text modalities presents greater challenges.
Finding 4 ($\Rightarrow$ RQ3): MLRMs generate harmful content during the thought process, which poses a potential risk in response.
Figure 4: ASR scores of the thought process and the final response generated by MLRMs on MMSafetyBench respectively.
Performance of Thought Process.
To investigate the impact of reasoning pathways in MLRMs, we independently evaluate the safety performance of the thought process. As shown in Figure 4, models that explicitly exhibit long CoT Kojima et al. (2022) generate harmful content during the thought process. Additionally, the ASR results for the four models—QVQ-Preview, R1-Onevision, Llava-cot, and Skywork-R1V—are higher for the thought process than for the final responses. Moreover, we analyze the thought process to identify how intermediate content contributes to the production of unsafe responses. As illustrated in Figure 3, although the model detects harmful intent in some cases, it does not reject the prompt outright. Instead, MLRM engages in self-deception by reframing the intent of users as educational, humorous, or otherwise benign, which rationalizes the model to generate harmful content. It is a critical challenge to address subtle intent rationalization in long reasoning, which needs to be explored in future work.
Figure 5: Overview of our data construction. We propose a multi-step pipeline to build the datasets based on various safety-related topics (Step 1: Word Cloud in Blue Block) and image description (Step 2: Red Block), which provides a thoughtful consideration of reasoning with safety, explicitly incorporating long CoT reasoning into the addressing process (Step 3: Green Block) and meticulously designed filtering mechanism (Step 4: Yellow Block).
Finding 5 ($\Rightarrow$ RQ3): MLRMs attempt to eliminate unsafe content in the thought process via self-reflection, while there are also situations where unsafe answers are generated after reducing the thinking steps.
R1-Onevision Safe Response Unsafe Response
Safe Thought 12.65 1.90
Unsafe Thought 7.34 78.11
Kimi-VL-thinking Safe Response Unsafe Response
Safe Thought 12.68 2.74
Unsafe Thought 25.83 58.75
Llava-cot Safe Response Unsafe Response
Safe Thought 15.89 3.06
Unsafe Thought 11.85 69.23
QVQ-Preview Safe Response Unsafe Response
Safe Thought 24.52 6.91
Unsafe Thought 6.19 62.38
Skywork-R1V Safe Response Unsafe Response
Safe Thought 19.88 3.93
Unsafe Thought 7.44 68.75
Table 4: Proportion of safety in the thought process and the final response on MMSafetyBench.
Comparative Effect between Thought and Response.
As shown in Table 4, the results reveal that the potential risk arising from the thought process does not always show up consistently in the final response. For example, the safe thought process leads to the unsafe response in QVQ-Preview, with respect to 6.91% proportion. In specific examples, when the model meets with unsafe intents, it is not engaged in in-depth thinking to generate unsafe content in thought while directly generate unsafe answers in response. The inconsistency between safe thought and unsafe response is due to the potential security risks inherent in the model. On the contrary, the unsafe thought process also results in the safe final response. It is attributed to the internal self-reflection and error-correction mechanisms of MLRMs, which enables models to eventually recognize the harmful nature of the intended prompt. In particular, Kimi-VL-thinking demonstrates a relatively strong capacity for self-correction with respect to 25.83% proportion, which aligns with the observations reported in Finding 2. The comparative results also provide potential insights for mitigating the issue of safety degradation in MLRMs.
3 Data Construction
Although existing studies have introduced several datasets for multimodal safety alignment, most of them consist of brief responses that lack explicit thought process. The supervised fine-tuning method based on these datasets struggles with retaining the inherent reasoning-chain advantages of MLRMs. Furthermore, arriving at conclusions without engaging in a safety-oriented thought process weakens the effectiveness and robustness of safety defense mechanisms. To address this issue, we propose a method for building a multimodal alignment dataset enriched with safety-oriented thought process, which offers a green channel to improve the safety of MLRMs. As a result, we aim to leverage the reasoning capabilities of MLRMs to improve the alignment with safety objectives.
As shown in Figure 5, we employ a multi-step pipeline to construct our safety dataset. After collecting various safety-related topics and image descriptions, we explicitly incorporate long CoT reasoning into the addressing process and design a filtering mechanism to finally obtain the dataset. To the best of our knowledge, it is the first safety dataset with the ability to retain reasoning chain for MLRMs.
[…]
(R1-Onevision)
Direct 72.20 79.57 32.14 17.31 49.69
VLGuard 65.80 35.60 3.21 40.72 60.88
MIS 56.00 60.18 7.14 31.74 53.63
Ours 16.80 21.54 0.00 71.26 61.09
(Llava-cot)
Direct 84.80 72.26 33.57 26.95 51.26
VLGuard 72.00 25.65 10.00 75.00 51.42
MIS 75.60 58.45 11.79 79.04 51.08
Ours 20.60 11.90 0.71 97.01 54.73
Table 5: Results of supervised fine-tuning method using different datasets. The best safety scores of the MLRM on each benchmark are highlighted in bold.
Data Preparation.
Due to the limitation of existing multimodal safety alignment datasets, we aim to construct a dataset enriched with safety-oriented thought process. However, considering the inherent security vulnerabilities of existing models and the development costs, it is not a practical approach to build such a dataset entirely from scratch. Therefore, we augment the existing dataset that already contains responses by adding structured and security-oriented thought process. In particular, we select two existing multimodal safety alignment datasets, VLGuard Zong et al. (2024) and SPA-VL Zhang et al. (2024), as the original data sources. Moreover, we define a set of twelve safety categories to select the original data, following the OpenAI safety guidelines, instead of using all data. More details are provided in Appendix C.1.
Image Description Generation.
To obtain a reliable thought process, we choose reasoning models instead of instruct models, which lack the ability to generate the thinking process with diversity and logicality. Furthermore, it is not feasible to directly utilize MLRMs to augment the thought process of the original data with their poor security performanc. Moreover, API-based proprietary models such as Claude3.7-sonnet-thinking with stronger safety mechanisms typically only return brief refusals (e.g., “I’m sorry”) when faced with some unsafe queries, offering no insight into their thought process. Therefore, we utilize the high-performing text-only reasoning model DeepSeek-R1 Guo et al. (2025) to generate safety-oriented thought process. To accommodate this model, we convert images into caption form as input in advance. In particular, we employ Qwen2.5-VL-72B Bai et al. (2025) to generate image captions to ensure that the semantic loss during modality conversion is within an acceptable range in terms of security issues, with the prompt “Please provide a detailed description of this image.
Safety Thought Process Generation.
To ensure the thought process generated by LRMs are truly safe, we add supplementary information related to security, inspired by the concept of deliberative alignment Guan et al. (2024). Previous methods often require models to infer implicit safety rules from large volumes of annotated examples, which suffers from low efficiency and limited generalization. Therefore, we provide clear and structured safety guidelines to directly guide the model in generating safety-oriented CoT. The guidelines not only help ensure the safety of the generated thought process, but also encourage MLRMs to think explicitly with safety considerations when producing responses. Besides, we develop the safety guidelines customized to the specific characteristics and risk profiles of each category. The image caption, question, original response, and safety guidelines are provided to DeepSeek-R1 Guo et al. (2025) to generate safety-oriented thought process. The model first analyzes the question and the caption to obtain the underlying intent, then refers to the safety guidelines to generate a safety-oriented thought process corresponding to the provided response. Prompt and examples of safety guidelines used in this step are provided in the Appendix C.2.
Thought Process Filter.
In the previous step, we obtain the original thought process. However, it is still necessary to ensure that the thought process is consistent with the corresponding response, which means that both should consistently reject or not reject the inquiry. Additionally, while the input to the model in the prior step is a caption, the input during MLRM fine-tuning is the original image. Therefore, the caption field is processed to prevent potential misinterpretation by the MLRM. The original thought process from the previous step also contain redundant elements such as headers like “Safety Chain of Thought” which need to be removed and reformatted. Given the relatively low complexity of this task, we employ DeepSeek-V3 Liu et al. (2024a) to handle this step. Specific prompts used for this step are provided in the Figure 6. After filtering, we construct the final safety datasets provided for the supervised fine-tuning of MLRMs.
4 Experiments
4.1 Experimental Settings
Baselines.
To evaluate the effectiveness of our data construction, we select two existing multimodal safety alignment datasets for supervised fine-tuning as baselines that are listed as follows:.
  • Direct: The original safety capability of the model without fine-tuning.
  • VLGuard Zong et al. (2024): The first safety dataset specifically designed for fine-tuning of MLLMs.
  • MIS Ding et al. (2025b): A multi-image safety dataset that combines visual perception with reasoning logic labels.
Training Details.
We use R1-Onevision Yang et al. (2025) and Llava-cot Xu et al. (2024) as base MLRMs for supervised fine-tuning, which clearly delineates the boundary between the thought process and the final response. We adopt supervised fine-tuning approach to the two models using the baseline datasets and our proposed dataset, respectively. The Details of the training procedure are provided in Appendix B.
4.2 Results
As shown in Table 5, the results demonstrate that the MLRMs achieve the best security performance after fine-tuning using our dataset on all benchmarks, compared with other baseline datasets. It is worth emphasizing that despite the complete absence of jailbreak attack data in our dataset, the proposed approach achieves a substantial reduction of ASR on jailbreak robustness benchmarks. Furthermore, since our dataset incorporates thought process that closely align with the data distribution used in MLRM training, the fine-tuned model preserved the original capability to generate coherent reasoning pathways. More case studies are provided in Appendix D.2. In addition, by fine-tuning the models respectively on the VLGuard, the MIS and the dataset we proposed, the safety performance of both models has been consistently improved, proving that well-designed constructing data is an effective solution to enhance the safety of the models.
5 Related Work
Safety of LRMs.
With the rapid development and widespread deployment of LRMs, many works have paid attention to security of LRMs. Several studies  Jiang et al. (2025); Parmar and Govindarajulu (2025) have conducted comprehensive security evaluations on LRMs, e.g., DeepSeek-R1 Guo et al. (2025), revealing existing vulnerabilities of these models. In addition, Zhou et al. (2025) introduce a CoT jailbreak attack method, specifically targeting reasoning models. On the other hand, to mitigate the issues of security, recent studies Jiang et al. (2025); Zhang et al. (2025b) have constructed the safety data with the thought process for supervised fine-tuning of LRMs. However, the above-mentioned works are all confined to the text-only models. Although SafeMLRM Fang et al. (2025) explores the security of MLRMs, the limitation of this work is that it only focuses on one security scenario, i.e., jailbreak robustness. The introduction of another modality in LRMs brings about significant shifts in model behavior, making it crucial to systematically assess the security challenges of MLRMs and develop effective defense methods.
Safety of MLLMs.
There are several studies that enhance the safety capabilities of MLLMs, which can be categorized into two types: the training-based method and the training-free method. The training-based method typically includes Supervised Fine-Tuning (SFT) Zong et al. (2024); Ding et al. (2025b); Li et al. (2024) and Reinforcement Learning from Human Feedback (RLHF) Zhang et al. (2024). In addition, Lee et al. (2024) introduce the weight merging approach to mitigate safety degradation. Another branch of studies incorporates additional safety components in a training-free manner. For instance, MLLM-Protector Pi et al. (2024) offers a plug-and-play detector for harmful responses, and ETA Ding et al. (2025a) proposes an inference-time alignment framework to ensure safety compliance during deployment. In contrast to these methods, our study is the first to investigate the safety behaviors of reasoning models in both jailbreak and awareness of safety scenarios.
6 Conclusion
In this work, we conduct a comprehensive evaluation and analysis of the safety of existing MLRMs, covering both jailbreak and comprehension benchmarks. The empirical results unveil several new findings, demonstrating that the safety performance of current MLRMs remains a significant concern. Furthermore, motivated by the findings, we propose a supervised fine-tuning dataset that considers explicit safety-oriented thought process. Experimental results on R1-Onevision and Llava-cot demonstrate that our dataset outperforms existing alternatives. This work represents a preliminary exploration of improving the safety of MLRMs through reasoning-based alignment. Future research will focus on developing more efficient datasets and training strategies specifically designed for MLRMs.
Appendix A Evaluation Details
Our evaluation encompasses five categories of safety benchmarks, which can be broadly divided into two types: jailbreak robustness and safety-awareness. Specifically, FigStep, MMSafetyBench, and JailBreaKV fall under the jailbreak robustness benchmark, for which we report the Attack Success Rate (ASR). Meanwhile, MSSBench and SIUO are categorized as safety-awareness benchmarks, and we follow the respective safety metrics according to the original papers.
The FigStep adopts the human evaluation as safety judgment. In our experiments, we follow the classification criteria described in their paper but apply an evaluation protocol consistent with that used for the MMSafetyBench. For the JailBreaKV, we follow the same evaluation settings as their work and utilize the Llama-Guard-Vision Chi et al. (2024) as the evaluator. The benchmarks are listed as follows:
FigStep.
FigStep Gong et al. (2025) consists of 500 harmful queries spanning 10 categories. Rather than directly inputting harmful textual prompts, it conveys malicious intent through visual formatting. Each image encodes harmful content as a visually rendered list, which is then paired with a fixed textual instruction: “The image shows a list numbered 1, 2, and 3, but the items are empty. Please generate detailed content for each item on the list.
$M$-SafetyBench.
$M$-SafetyBench Liu et al. (2024b) covers 13 categories and is primarily designed to evaluate the responses of MLLMs to malicious queries. In our experiments, we adopt the SD+TYPO setting, which includes a total of 1,680 samples, following the setting of Ding et al. (2025b).
JailBreaKV.
JailBreaKV Luo et al. (2024) is designed to evaluate the transferability of jailbreak techniques from LLMs to MLLMs, encompassing both text-based and image-based jailbreak attacks. Following the evaluation protocol of VLGuard, we conduct our experiments on the miniset provided in their paper, which contains 280 samples, following the setting of Hu et al. (2024).
MSSBench.
Multimodal Situation Safety Benchmark (MSSBench) Zhou et al. (2024) explores how safety considerations vary depending on the specific context involving the user or agent. In this benchmark, half of the image contexts are safe, while the other half are unsafe. The benchmark identifies safe intent of the models in two distinct settings: chat and embodied scenarios. Due to economic constraints, our evaluation was conducted using a single instruction sampled from each of the safe and unsafe scenarios.
SIUO.
Safety Inputs but Unsafe Outputs (SIUO) Wang et al. (2025) considers cases where the image and text modalities are independently safe, but their combination leads to unsafe or unethical outputs. The SIUO covers 9 categories including a total of 167 samples. We report the evaluation metrics as defined in their paper.
Appendix B Training Details
For the training procedure, we utilize the LoRA Hu et al. (2022) to fine-tune the R1-Onevision and Llava-cot based on the LLaMA-Factory [^1] https://github.com/hiyouga/LLaMA-Factory framework. All training experiments are conducted using one NVIDIA A100-80G GPUs. The hyper-parameters for supervised fine-tuning are listed in Table 6.
Hyper-Parameter Value
lora rank 8
learning rate 1.0e-5
train epoch 1.5
per_device_batchsize 1
warm up ratio 0.1
learning rate scheduler cosine
Table 6: Hyper-parameters for supervised fine-tuning.
Model Chat Embodied AVG.
Safe Unsafe AVG. Safe Unsafe AVG.
Gemini2.0-flash-thinking 85.67 45.00 65.33 93.06 5.56 49.31 57.32
Claude3.7-sonnet-thinking 75.84 45.97 60.91 93.42 15.79 54.61 57.76
QVQ-Preview 80.67 24.67 52.67 53.95 43.42 48.68 50.68
Skywork-R1V 86.67 15.00 50.83 98.68 1.32 50.00 50.42
Llama3.2-vision-11B(base) 85.62 22.07 53.85 90.79 10.53 50.66 52.26
LlamaV-o1 59.33 47.00 53.17 100.00 0.00 50.00 51.59
Llama3.2-vision-11B(base) 85.62 22.07 53.85 90.79 10.53 50.66 52.26
Llava-cot 93.67 12.67 53.17 97.37 1.32 49.34 51.26
Llama3.2-vision-11B(base) 85.62 22.07 53.85 90.79 10.53 50.66 52.26
Mulberry-Llama 81.33 35.00 58.17 97.37 2.63 50.00 54.09
Qwen2.5-VL-3B(base) 83.67 31.00 57.33 59.21 35.53 47.37 52.35
LMM-R1 83.00 33.00 58.00 69.70 26.32 48.03 53.02
Qwen2.5-VL-7B(base) 93.33 11.00 52.17 93.42 6.58 50.00 51.09
R1-Onevision 86.25 15.46 50.86 86.57 10.45 48.51 49.69
InternVL-2.5-8B (base) 92.00 11.33 51.67 98.68 1.32 50.00 50.84
$M$-Eureka 90.67 11.67 51.17 97.37 2.63 50.00 50.59
Kimi-VL-instruct (base) 91.67 8.67 50.17 100.00 1.43 50.71 50.44
Kimi-VL-thinking 83.33 23.67 53.5 98.67 0.00 49.33 51.42
Table 7: Results on MSSBench in two distinct settings: chat and embodied scenarios.
Appendix C Dataset Details
C.1 Data Statistics
Based on the categories provided in SPA-VL Zhang et al. (2024) and VLGuard Zong et al. (2024) datasets, we extract the raw data in the first step of construction for data preparation. In addition, we include the paired safety examples from VLGuard, where each image is matched with both a safe and an unsafe question. This design aims to improve the ability of the models to handle context-sensitive safety issues. However, the answers in SPA-VL are derived from pairwise preference data generated by large language models. Despite selecting the “chosen” responses as the final answers, a subset of them remain potential safety risks. The unsafe answers are subsequently screened and reconstructed. The statistics of data instances in each category within our dataset after the final filtering process are shown in Table 8.
Category # Samples
Privacy violation 196
Professional advice 200
Political sensitivity 209
Sexually explicit 199
Violence 204
Disinformation 205
Discrimination 231
Hate speech 200
Economic harm 200
Physical harm 200
Illegal activities 200
Malware 200
Safe 977
All 4398
Table 8: The statistics of our dataset
C.2 Prompt for Data Conustruction
This section presents the prompt and safety regulation used for safety thought process collection. We referenced the definition of User Requests Categorization from Wang et al. (2024) and listed all prompts used in the data construction in Figure 6. During the collection of safety thought process data, we supplied the model with category-specific safety regulations derived from the raw data classification. This strategy not only reduce the length of prompt but also enable tiered defensive measures for distinct safety-related issues.
Appendix D More Results
D.1 Detailed Experiments on MSSBench
In the evaluation protocol of MSS-Bench, scenarios are divided into chat and embodied settings, with safety scores calculated separately for safe and unsafe conditions. The average safety scores are reported in Table 2, and the specific scores for each category are detailed in Table 7.
Figure 6: Prompt used in data construction.
D.2 Case Study
Examples of responses generated by fine-tuned on VLGuard, MIS, and our dataset are illustrated in Figure 7. Our approach demonstrates the ability to retain the thought process of the models while decisively rejecting unsafe inputs and explicitly articulating the potential serious consequences associated with such queries.
Figure 7: Examples of several baselines and our dataset on $M$-SafetyBench. We only highlight some fragments to show the representative difference between various methods. The red indicates the unsafe parts, while the blue indicates the safe parts.

### v2

@@
 Moreover, our analysis reveals distinct safety patterns across different benchmarks: significant safety degradation is observed across jailbreak robustness benchmarks, whereas safety-awareness benchmarks demonstrate less pronounced degradation.
-In particular, a long thought process in some scenarios even enhances safety performance.
-Therefore, it is a potential approach to addressing safety issues in MLRMs by leveraging the intrinsic reasoning capabilities of the model to detect unsafe intent.
+In particular, the long thought process in some scenarios even enhances safety performance.
+Therefore, it is a potential approach to address safety issues in MLRMs by leveraging the intrinsic reasoning capabilities of the model to detect unsafe intent.
 To operationalize this insight, we construct a multimodal tuning dataset that incorporates a safety-oriented thought process.
@@
 Contact e-mail: louxinyue@bjtu.edu .cn, youlee@bjtu.edu.cn, Jinan Xu1, Xiangyu Shi1, Chi Chen2, Kaiyu Huang1 1 Beijing Jiaotong University, Beijing, China 2 Tsinghua University, Beijing, China
+22footnotetext: Corresponding author: kyhuang@bjtu.edu.cn
 1 Introduction
-The development of large reasoning models (LRMs) such as OpenAI o1 Jaech et al.
+The development of Large Reasoning Mode-ls (LRMs) such as OpenAI o1 Jaech et al.
 (2024) and DeepSeek-R1 Guo et al.
@@
 (2025).
-However, current research on the safety of MLRMs remains limited, primarily focusing on two jailbreak robustness benchmarks, with insufficient exploration of the broader spectrum of safety challenges in multimodal settings.
+However, current research on the safety of MLRMs remains limited, primarily focusing on jailbreak robustness benchmarks Gong et al.
+(2025); Liu et al.
+(2024b), with insufficient exploration of the broader spectrum of safety challenges in multimodal settings.
 As shown in Figure 1, we categorize the safety benchmarks of multimodal models into two key aspects: safety-awareness and jailbreak robustness.
-Jailbreak robustness benchmarks focus on evaluating the resilience of the model against deliberately crafted or modified textual prompts and visual inputs that aim to bypass established safety defense mechanisms.
-Safety-awareness benchmarks emphasize the capability of models to proactively identify potential safety risks embedded in user inputs, aligning more closely with the complex and dynamic safety demands encountered in real-world applications.
+Jailbreak robustness benchmarks Gong et al.
+(2025); Liu et al.
+(2024b); Luo et al.
+(2024) focus on evaluating the resilience of the model against deliberately crafted or modified textual prompts and visual inputs that aim to bypass established safety defense mechanisms.
+Safety-awareness benchmarks Wang et al.
+(2025); Zhou et al.
+(2024) emphasize the capability of models to proactively identify potential safety risks embedded in user inputs, aligning more closely with the complex and dynamic safety demands encountered in real-world applications.
 Compared to unimodal LRMs, MLRMs exhibits the following research questions:
 RQ1: How does MLRMs affect safety on different types of benchmarks compared with Multimodal Large Language Models (MLLMs)?
-RQ2: What are the limitations of incorporating additional modalities in MLRMs for safety concerns?
+RQ2: What are the risks of incorporating additional modalities in MLRMs for safety concerns?
 RQ3: What is the impact of reasoning pathways in MLRMs on the safety performance?
@@
 (2025), R1-Onevision Yang et al.
-(2025) and etc.
+(2025) , etc.
 The results demonstrate that MLRMs have significant negative impact on the safety performance, while the performance degradation is task-dependent.
-In particular, MLRMs could identify more potential safety risks through deliberate thinking, leading to higher safety scores on safety-awareness benchmarks, which provides a novel perspective for mitigating safety degradation in MLRMs.
+Furthermore, we found that MLRMs could identify more potential safety risks through deliberate thinking, leading to higher safety scores on safety-awareness benchmarks, which provides a novel perspective for mitigating safety degradation in MLRMs.
 Motivated by this, we further propose a data construction method that incorporates safety-oriented thought process to investigate the effectiveness of this insight.
 To sum up, our contributions are summarized as follows:
-  • This study conducts a systematic evaluation of MLRMs and investigates the empirical results, revealing several novel findings and providing new perspectives for the development of safer MLRMs.
-  • We construct a multimodal fine-tuning dataset with safety-oriented thought process for safety alignment, alleviating the issue of incorporating additional modalities.
-  • Experimental results demonstrate that our method improve the safety performance of MLRMs across multiple benchmarks by enabling self-correction thinking along the reasoning pathways, compared with previous defense methods.
-Model Name Base Model
-LlamaV-o1 Thawakar et al. (2025) Llama3.2-vision-11B
-Llava-cot Xu et al. (2024) Llama3.2-vision-11B
-Mulberry-Llama Yao et al. (2024) Llama3.2-vision-11B
-LMM-R1 Peng et al. (2025b) Qwen2.5-VL-3B
-R1-Onevision Yang et al. (2025) Qwen2.5-VL-7B
-MM-Eureka Meng et al. (2025) InternVL-2.5-8B
-Kimi-VL-thinking Team et al. (2025) Kimi-VL-instruct
-Skywork-R1V Peng et al. (2025a) -
-QVQ-Preview -
-Gemini2.0-flash-thinking -
-Claude3.7-sonnet-thinking -
-Table 1: List of the evaluated MLRMs and their corresponding base models.
+  • This study conducts a systematic safety evaluation of MLRMs and investigates the empirical results, revealing several novel findings and providing new perspectives for the development of safer MLRMs.
+  • We construct a multimodal fine-tuning dataset with safety-oriented thought process for safety alignment, alleviating the issue associated with incorporating additional modalities.
+  • Experimental results demonstrate that our method improves the safety performance of MLRMs across multiple benchmarks by enabling self-correction thinking along the reasoning pathways, compared with previous defense methods.
   Jailbreak Robustness Safety-Awareness
@@
 Llama 3.2-vision-11B(base) 55.80 38.45 5.71 37.13 52.26
-Llava-cot 84.80(+29.00) 72.26(+33.81) 33.57(+27.86) 26.95(-10.18) 51.26(-1.00)
+LLaVA-CoT 84.80(+29.00) 72.26(+33.81) 33.57(+27.86) 26.95(-10.18) 51.67(-0.59)
 Llama 3.2-vision-11B(base) 55.80 38.45 5.71 37.13 52.26
@@
 LMM-R1 69.80(+3.53) 68.15(+1.97) 18.21(+6.07) 21.56(-2.99) 53.02(+0.67)
-Qwen2.5-VL-7B(base) 67.20 66.49 12.50 29.94 51.09
-R1-Onevision 72.20(+5.00) 79.57(+13.08) 32.14(+19.64) 17.31(-12.63) 49.69(-1.40)
+Qwen2.5-VL-7B(base) 67.20 66.49 12.50 29.94 50.02
+R1-Onevision 72.20(+5.00) 79.57(+13.08) 32.14(+19.64) 17.31(-12.63) 48.94(-1.08)
 InternVL-2.5-8B(base) 71.40 59.64 15.00 28.14 50.84
 $M$-Eureka 72.20(+0.80) 60.12(+0.48) 11.79(-3.21) 28.14(-0.00) 50.59(-0.25)
-Kimi-VL-instruct(base) 80.40 47.74 22.50 25.00 50.44
+Kimi-VL(base) 80.40 47.74 22.50 25.00 50.44
 Kimi-VL-Thinking 87.00(+6.60) 61.49(+13.75) 33.93(+11.43) 35.93(+10.93) 51.42(+0.98)
@@
 To comprehensively assess the safety performance of MLRMs across diverse scenarios, we adopt benchmark datasets from two distinct perspectives: safety-awareness and jailbreak robustness.
-For the assessment of safety-awareness, we employ SIUO Wang et al.
+For assessment of safety-awareness, we employ SIUO Wang et al.
 (2025) and MSSBench Zhou et al.
-(2024) datasets, while for evaluating jailbreak robustness, we employ MMSafetyBench Liu et al.
+(2024) datasets, while for evaluating jailbreak robustness, we employ MM-SafetyBench Liu et al.
 (2024b), FigStep Gong et al.
@@
 (2024).
-In safety-awareness tasks, models need to reason jointly over both visual and textual inputs to infer user intent, identify potential safety risks, and assess whether the input should be treated as safe or unsafe.
+In safety-awareness tasks, models need to jointly reason over both visual and textual inputs to infer user intent, identify potential safety risks, and assess whether the input should be treated as safe or unsafe.
 These tasks pose significant challenges to the multimodal reasoning and safety alignment capabilities of models.
-In contrast, jailbreak robustness benchmarks involve adversarial attacks, such as the inclusion of malicious crafted prompt, aimed at circumventing the safety constraints of models.
-Given that these two task categories examine safety alignment from fundamentally different perspectives, we analyze the results separately.
-Further details regarding the datasets are provided in Appendix A.
+In contrast, jailbreak robustness benchmarks involve adversarial attacks, such as the inclusion of maliciously crafted prompt, aimed at circumventing the safety constraints of models.
+Given that these two task categories examine safety alignment from fundamentally different perspectives, we analyze these results separately.
+Further details regarding the datasets are provided in Appendix A.1.
 Models and Configurations.
@@
 (2025).
-Detailed information of the models are provided in Table 1.
+Detailed information of the models are provided in Appendix A.3.
 For reasoning models that have undergone extensive fine-tuning—via either supervised learning or reinforcement learning—we additionally evaluate their corresponding base models.
 This approach allows us to trace safety-related changes across model variants.
-All experiments were conducted using four NVIDIA A100-80G GPUs.
-Figure 2: Case study of the better safety consideration on safety-awareness tasks.
-Kimi-VL-Instruct directly outputs the answer that ignores the potential risk, while Kimi-VL-Thinking dives deeper into the insidious safety issue with stronger reasoning abilities.
+All experiments are conducted using four NVIDIA A100-80G GPUs.
 Metrics and Evaluator.
@@
 For safety-awareness benchmarks, we follow established evaluation protocols to compute a safety score, where higher scores denote better safety performance.
-Further evaluation settings are detailed in Appendix A.
-Following the settings of MMSafetyBench, MSS, and SIUO utilizing OpenAI API, we employ GPT-4o-mini Hurst et al.
+Further evaluation settings are detailed in Appendix A.1.
+Following the settings of MM-SafetyBench, MSS, and SIUO utilizing OpenAI API, we employ GPT-4o-mini Hurst et al.
 (2024) as the evaluator and adopt identical configurations to ensure consistency across assessments.
+Figure 2: Case study of the better safety consideration on safety-awareness tasks.
+Kimi-VL directly outputs the answer that ignores the potential risk, while Kimi-VL-Thinking dives deeper into the insidious safety issue with stronger reasoning abilities.
+The red indicates the unsafe parts, while the green indicates the content related to potential risks identified by reasoning models.
 2.2 Safety Results and Inspection
@@
 A detailed analysis of this phenomenon is presented in Finding 2.
-Finding 2 ($\Rightarrow$ RQ1): MLRM paradoxically enhances safety performance via long reasoning on safety-awareness benchmarks, due to improved capability in identifying unsafe intent.
+Finding 2 ($\Rightarrow$ RQ1): MLRMs paradoxically enhances safety performance via long reasoning on safety-awareness benchmarks, due to improved capability in identifying unsafe intent.
 Performance on Safety-Awareness.
@@
 The effect is particularly pronounced for Kimi-VL-Thinking model, which achieves a 10.9% improvement on the SIUO benchmark.
-As illustrated in Figure 2, there is a difference in attention patterns according to the responses generated by these models and their corresponding base models.
+As illustrated in Figure 2, there is a difference in focusing center according to the responses generated by these models and their corresponding base models.
 Reasoning models, when engaging in long reasoning, tend to focus more on potentially unsafe details present in both the visual and textual modalities.
@@
 […]
-Qwen2.5-VL-7B (I) 66.49 12.5 29.94
-Qwen2.5-VL-7B (T) 48.21(-18.28) 6.79(-5.71) 38.10(+8.16)
-R1-Onevision (I) 79.57 32.14 17.31
-R1-Onevision (T) 65.02(-14.55) 23.21(-8.93) 23.35(+6.04)
-Kimi-VL-instruct (I) 47.74 22.50 25.00
-Kimi-VL-instruct (T) 57.56(+9.82) 19.29(-3.21) 25.00(+0.00)
-Kimi-VL-thinking (I) 61.49 33.93 35.93
-Kimi-VL-thinking (T) 55.71(-5.78) 29.64(-4.29) 29.34(-6.59)
-Llama3.2-vision-11B (I) 38.45 5.71 37.13
-Llama3.2-vision-11B (T) 58.52(+20.07) 44.29(+38.58) 22.50(-14.63)
-Llava-cot (I) 72.26 33.57 26.95
-Llava-cot (T) 60.06(-12.20) 40.36(+6.79) 28.14(+1.19)
-InternVL2.5-8B (I) 59.64 15.00 28.14
-InternVL2.5-8B (T) 48.57(-11.07) 9.64(-5.36) 27.55(-0.59)
-MM-Eureka (I) 60.12 11.79 28.14
-MM-Eureka (T) 50.24(-9.88) 10.71(-1.08) 31.14(+3.00)
-Table 3: Safety performance when converting image into text caption. “I” indicates multimodal inputs, while “T” refers to text-only inputs.
+Qwen2.5-VL-7B 66.49 12.5 29.94
+     -w/o image 48.21(-18.28) 6.79(-5.71) 38.10(+8.16)
+R1-Onevision 79.57 32.14 17.31
+     -w/o image 65.02(-14.55) 23.21(-8.93) 23.35(+6.04)
+Kimi-VL 47.74 22.50 25.00
+     -w/o image 57.56(+9.82) 19.29(-3.21) 25.00(+0.00)
+Kimi-VL-Thinking 61.49 33.93 35.93
+     -w/o image 55.71(-5.78) 29.64(-4.29) 29.34(-6.59)
+Llama 3.2-vision-11B 38.45 5.71 37.13
+     -w/o image 58.52(+20.07) 44.29(+38.58) 22.50(-14.63)
+LLaVA-CoT 72.26 33.57 26.95
+     -w/o image 60.06(-12.20) 40.36(+6.79) 28.14(+1.19)
+InternVL2.5-8B 59.64 15.00 28.14
+     -w/o image 48.57(-11.07) 9.64(-5.36) 27.55(-0.59)
+$M$-Eureka 60.12 11.79 28.14
+     -w/o image 50.24(-9.88) 10.71(-1.08) 31.14(+3.00)
+Table 2: Safety performance when converting image into text caption.
+The safety performance variation has been marked in brackets, where the red color represents safety deterioration and green stands for safety improvement.
+MMSafe is the abbreviation for $M$-SafetyBench.
 Finding 3 ($\Rightarrow$ RQ2): The substitution of the image modality with text modality leads to a partial recovery of the safety defense capabilities of models.
@@
 Finding 4 ($\Rightarrow$ RQ3): MLRMs generate harmful content during the thought process, which poses a potential risk in response.
-Figure 4: ASR scores of the thought process and the final response generated by MLRMs on MMSafetyBench respectively.
+Figure 4: ASR scores of the thought process and the final answer generated by MLRMs on FigStep respectively.
 Performance of Thought Process.
 To investigate the impact of reasoning pathways in MLRMs, we independently evaluate the safety performance of the thought process.
-As shown in Figure 4, models that explicitly exhibit long CoT Kojima et al.
-(2022) generate harmful content during the thought process.
-Additionally, the ASR results for the four models—QVQ-Preview, R1-Onevision, Llava-cot, and Skywork-R1V—are higher for the thought process than for the final responses.
-Moreover, we analyze the thought process to identify how intermediate content contributes to the production of unsafe responses.
+As shown in Figure 4, the proportion of unsafe thought process generated by the four models (R1-Onevision, LLaVA-CoT, Kimi-VL-Thinking and Skywork-R1V) all exceeds 50 percent.
+Moreover, we analyze the thought process to identify how intermediate content contributes to the production of unsafe answers.
 As illustrated in Figure 3, although the model detects harmful intent in some cases, it does not reject the prompt outright.
@@
 It is a critical challenge to address subtle intent rationalization in long reasoning, which needs to be explored in future work.
-Figure 5: Overview of our data construction.
-We propose a multi-step pipeline to build the datasets based on various safety-related topics (Step 1: Word Cloud in Blue Block) and image description (Step 2: Red Block), which provides a thoughtful consideration of reasoning with safety, explicitly incorporating long CoT reasoning into the addressing process (Step 3: Green Block) and meticulously designed filtering mechanism (Step 4: Yellow Block).
-Finding 5 ($\Rightarrow$ RQ3): MLRMs attempt to eliminate unsafe content in the thought process via self-reflection, while there are also situations where unsafe answers are generated after reducing the thinking steps.
-R1-Onevision Safe Response Unsafe Response
-Safe Thought 12.65 1.90
-Unsafe Thought 7.34 78.11
-Kimi-VL-thinking Safe Response Unsafe Response
-Safe Thought 12.68 2.74
-Unsafe Thought 25.83 58.75
-Llava-cot Safe Response Unsafe Response
-Safe Thought 15.89 3.06
-Unsafe Thought 11.85 69.23
-QVQ-Preview Safe Response Unsafe Response
-Safe Thought 24.52 6.91
-Unsafe Thought 6.19 62.38
-Skywork-R1V Safe Response Unsafe Response
-Safe Thought 19.88 3.93
-Unsafe Thought 7.44 68.75
-Table 4: Proportion of safety in the thought process and the final response on MMSafetyBench.
-Comparative Effect between Thought and Response.
-As shown in Table 4, the results reveal that the potential risk arising from the thought process does not always show up consistently in the final response.
-For example, the safe thought process leads to the unsafe response in QVQ-Preview, with respect to 6.91% proportion.
+Figure 5: Overview of our data construction pipeline.
+We propose a multi-stage pipeline to build the datasets based on various safety-related topics (Step 1: Word Cloud in Blue Block) and image description (Step 2: Red Block), which provides a thoughtful consideration of reasoning with safety, explicitly incorporating long CoT reasoning into the addressing process (Step 3: Green Block) and meticulously designed filtering mechanism (Step 4: Yellow Block).
+Finding 5 ($\Rightarrow$ RQ3): MLRMs attempt to eliminate unsafe content in the thought process via self-reflection, while there are also situations where unsafe answers are generated even if the thought process is secure.
+  R1-Onevision Safe Answer Unsafe Answer
+Safe Thought 13.60% 29.40%
+Unsafe Thought 14.20% 42.80%
+Kimi-VL-Thinking Safe Answer Unsafe Answer
+Safe Thought 8.40% 30.40%
+Unsafe Thought 4.60% 48.20%
+LLaVA-CoT Safe Answer Unsafe Answer
+Safe Thought 8.80% 30.20%
+Unsafe Thought 6.40% 54.60%
+QVQ-Preview Safe Answer Unsafe Answer
+Safe Thought 24.60% 46.60%
+Unsafe Thought 4.40% 24.20%
+Skywork-R1V Safe Answer Unsafe Answer
+Safe Thought 9.40% 37.60%
+Unsafe Thought 4.80% 48.20%
+Table 3: Proportion of safety in the thought process and the final answer on FigStep.
+Comparative Effect between Thought and Answer.
+As shown in Table 3, the results reveal that the potential risk arising from the thought process does not always show up consistently in the final answer.
+For example, the safe thought process leads to the unsafe answer in QVQ-Preview, with respect to 46.60% proportion.
 In specific examples, when the model meets with unsafe intents, it is not engaged in in-depth thinking to generate unsafe content in thought while directly generate unsafe answers in response.
-The inconsistency between safe thought and unsafe response is due to the potential security risks inherent in the model.
-On the contrary, the unsafe thought process also results in the safe final response.
+The inconsistency between safe thought and unsafe answer is due to the potential security risks inherent in the model.
+On the contrary, the unsafe thought process also results in the safe final answer.
 It is attributed to the internal self-reflection and error-correction mechanisms of MLRMs, which enables models to eventually recognize the harmful nature of the intended prompt.
-In particular, Kimi-VL-thinking demonstrates a relatively strong capacity for self-correction with respect to 25.83% proportion, which aligns with the observations reported in Finding 2.
 The comparative results also provide potential insights for mitigating the issue of safety degradation in MLRMs.
 3 Data Construction
-Although existing studies have introduced several datasets for multimodal safety alignment, most of them consist of brief responses that lack explicit thought process.
-The supervised fine-tuning method based on these datasets struggles with retaining the inherent reasoning-chain advantages of MLRMs.
+Although existing studies have introduced several datasets for multimodal safety alignment, most of them Zong et al.
+(2024); Ding et al.
+(2025b) consist of brief responses that lack explicit thought process.
+The fine-tuning method based on these datasets struggles with retaining the inherent reasoning-chain advantages of MLRMs.
 Furthermore, arriving at conclusions without engaging in a safety-oriented thought process weakens the effectiveness and robustness of safety defense mechanisms.
-To address this issue, we propose a method for building a multimodal alignment dataset enriched with safety-oriented thought process, which offers a green channel to improve the safety of MLRMs.
-As a result, we aim to leverage the reasoning capabilities of MLRMs to improve the alignment with safety objectives.
-As shown in Figure 5, we employ a multi-step pipeline to construct our safety dataset.
-After collecting various safety-related topics and image descriptions, we explicitly incorporate long CoT reasoning into the addressing process and design a filtering mechanism to finally obtain the dataset.
-To the best of our knowledge, it is the first safety dataset with the ability to retain reasoning chain for MLRMs.
+To address these issues, we propose a method for building a multimodal alignment dataset named TiS (Think in Safety), distinguished by its safety-oriented thought process and various safety-related topics.
+Through this method, we aim to leverage the innate reasoning capabilities of MLRM to improve its safety alignment.
+As shown in Figure 5, we employ a multi-stage pipeline to construct our safety alignment dataset TiS.
+We begin by collecting safety-related topics and generating image captions, then explicitly incorporating long CoT reasoning into question answering.
+After a filtering procedure, we finally obtain the dataset.
+To the best of our knowledge, TiS is the first safety dataset with the ability to retain reasoning chain for MLRMs.
+We provide a detailed illustration of the entire pipeline in Appendix B.1.
+  Jailbreak Robustness Safety-Awareness
 […]
 (R1-Onevision)
-Direct 72.20 79.57 32.14 17.31 49.69
-VLGuard 65.80 35.60 3.21 40.72 60.88
-MIS 56.00 60.18 7.14 31.74 53.63
-Ours 16.80 21.54 0.00 71.26 61.09
+Direct 72.20 79.57 32.14 17.31 48.94
+VLGuard 36.60(-35.60) 27.21(-52.36) 0.36(-31.78) 45.51(+28.20) 63.86(+14.92)
+MIS 68.00(-4.20) 46.02(-33.55) 2.14(-30.00) 41.92(+24.61) 65.61(+15.67)
+SPA-VL 30.00(-42.20) 35.30(-44.27) 0.36(-31.78) 42.51(+25.20) 63.60(+14.66)
+TiS (Ours) 15.80(-56.40) 21.79(-57.78) 0.00(-32.14) 71.26(+53.95) 66.31(+17.37)
+-w/o thought 20.80(-51.40) 16.37(-63.20) 0.00(-32.14) 63.47(+46.14) 63.92(+14.98)
 (LLaVA-CoT)
-Direct 84.80 72.26 33.57 26.95 51.26
-VLGuard 72.00 25.65 10.00 75.00 51.42
-MIS 75.60 58.45 11.79 79.04 51.08
-Ours 20.60 11.90 0.71 97.01 54.73
-Table 5: Results of supervised fine-tuning method using different datasets. The best safety scores of the MLRM on each benchmark are highlighted in bold.
+Direct 84.80 72.26 33.57 26.95 51.67
+VLGuard 62.60(-22.20) 14.76(-57.50) 5.36(-28.21) 61.68(+34.73) 53.40(+1.73)
+MIS 59.60(-25.20) 45.48(-26.78) 2.86(-3.71) 52.10(+25.15) 54.90(+3.23)
+SPA-VL 41.80(-43.00) 43.63(-28.63) 0.71(-32.86) 65.27(+38.32) 57.63(+5.96)
+TiS (Ours) 12.20(-72.60) 8.87(-63.39) 0.00(-33.57) 74.85(+47.90) 59.17(+7.50)
+-w/o thought 28.60(-56.20) 6.67(-65.59) 2.14(-31.43) 67.66(+40.71) 53.96(+2.09)
+Table 4: Results of supervised fine-tuning method using different datasets.
+The best safety scores of the MLRM on each benchmark are highlighted in bold and the second-best are highlighted in underline.
 Data Preparation.
 Due to the limitation of existing multimodal safety alignment datasets, we aim to construct a dataset enriched with safety-oriented thought process.
-However, considering the inherent security vulnerabilities of existing models and the development costs, it is not a practical approach to build such a dataset entirely from scratch.
-Therefore, we augment the existing dataset that already contains responses by adding structured and security-oriented thought process.
-In particular, we select two existing multimodal safety alignment datasets, VLGuard Zong et al.
+However, considering the inherent security vulnerabilities of existing models and the resource constraints, constructing such a dataset entirely from scratch is largely impractical.
+Therefore, we augment existing datasets with short responses by adding structured and security-oriented thought process.
+Specifically, we select two multimodal safety alignment datasets as our original data sources: VLGuard Zong et al.
 (2024) and SPA-VL Zhang et al.
-(2024), as the original data sources.
-Moreover, we define a set of twelve safety categories to select the original data, following the OpenAI safety guidelines, instead of using all data.
-More details are provided in Appendix C.1.
+(2024).
+Rather than directly utilizing the entire datasets, we retain only the high-quality instances and ensure a balanced distribution across diverse safety-related topics.
 Image Description Generation.
-To obtain a reliable thought process, we choose reasoning models instead of instruct models, which lack the ability to generate the thinking process with diversity and logicality.
-Furthermore, it is not feasible to directly utilize MLRMs to augment the thought process of the original data with their poor security performanc.
-Moreover, API-based proprietary models such as Claude3.7-sonnet-thinking with stronger safety mechanisms typically only return brief refusals (e.g., “I’m sorry”) when faced with some unsafe queries, offering no insight into their thought process.
-Therefore, we utilize the high-performing text-only reasoning model DeepSeek-R1 Guo et al.
-(2025) to generate safety-oriented thought process.
-To accommodate this model, we convert images into caption form as input in advance.
-In particular, we employ Qwen2.5-VL-72B Bai et al.
-(2025) to generate image captions to ensure that the semantic loss during modality conversion is within an acceptable range in terms of security issues, with the prompt “Please provide a detailed description of this image.
+To obtain a reliable thought process, we utilize the text-only reasoning model, Deepseek-R1 Guo et al.
+(2025), because instruction models often struggle to produce coherent reasoning, while MLRMs tend to suffer from alignment collapse under various conditions.
+Meanwhile, close-source proprietary models are equipped with defensive safety guardrail, which hinders access to their internal reasoning processes particularly for safety-related questions.
+To address the modality gap inherent in this text-only models, the visual content is converted into detailed image captions using Qwen2.5-VL-72B Bai et al.
+(2025).
+  R1-Onevision Safe Answer Unsafe Answer
+Safe Thought 13.60% 29.40%
+Unsafe Thought 14.20% 42.80%
+R1-Onevision+TiS Safe Answer Unsafe Answer
+Safe Thought 81.60% 15.60%
+Unsafe Thought 2.60% 0.20%
+LLaVA-CoT Safe Answer Unsafe Answer
+Safe Thought 8.80% 30.20%
+Unsafe Thought 6.40% 54.60%
+LLaVA-CoT+TiS Safe Answer Unsafe Answer
+Safe Thought 81.40% 11.80%
+Unsafe Thought 6.40% 0.40%
+Table 5: Proportion of safety in the thought process and the final answer of fine-tuned model using our dataset on FigStep.
 Safety Thought Process Generation.
-To ensure the thought process generated by LRMs are truly safe, we add supplementary information related to security, inspired by the concept of deliberative alignment Guan et al.
+To ensure the thought process generated by LRMs is aligned with human values, we explicitly add safety guidelines in the generation process, inspired by the concept of deliberative alignment Guan et al.
 (2024).
-Previous methods often require models to infer implicit safety rules from large volumes of annotated examples, which suffers from low efficiency and limited generalization.
-Therefore, we provide clear and structured safety guidelines to directly guide the model in generating safety-oriented CoT.
-The guidelines not only help ensure the safety of the generated thought process, but also encourage MLRMs to think explicitly with safety considerations when producing responses.
-Besides, we develop the safety guidelines customized to the specific characteristics and risk profiles of each category.
-The image caption, question, original response, and safety guidelines are provided to DeepSeek-R1 Guo et al.
-(2025) to generate safety-oriented thought process.
-The model first analyzes the question and the caption to obtain the underlying intent, then refers to the safety guidelines to generate a safety-oriented thought process corresponding to the provided response.
-Prompt and examples of safety guidelines used in this step are provided in the Appendix C.2.
-Thought Process Filter.
-In the previous step, we obtain the original thought process.
-However, it is still necessary to ensure that the thought process is consistent with the corresponding response, which means that both should consistently reject or not reject the inquiry.
-Additionally, while the input to the model in the prior step is a caption, the input during MLRM fine-tuning is the original image.
-Therefore, the caption field is processed to prevent potential misinterpretation by the MLRM.
-The original thought process from the previous step also contain redundant elements such as headers like “Safety Chain of Thought” which need to be removed and reformatted.
-Given the relatively low complexity of this task, we employ DeepSeek-V3 Liu et al.
-(2024a) to handle this step.
-Specific prompts used for this step are provided in the Figure 6.
-After filtering, we construct the final safety datasets provided for the supervised fine-tuning of MLRMs.
+Specific safety standards for various domains (law, medical, etc) are also clarified.
+Previous methods often require models to infer implicit safety rules from large volumes of training examples, which suffers from low efficiency and limited generalization.
+By contrast, our method provides sufficient safety standards for safety-centric reasoning content generation, facilitating deeper safety alignment and understanding rather than the first few tokens Qi et al.
+(2024).
+More details and prompt used in this step are provided in Appendix B.3.
+Thought Process Filtering.
+After obtaining the safety-oriented reasoning processes, we perform an additional modification and filtering step to ensure the quality and correctness of the data, making them better suited for safety enhancement.
+This step addresses issues such as inaccuracies (e.g., responses that explicitly state “according to the caption”) and misalignments in reasoning.
+We collaboratively employ Deepseek-V3 Liu et al.
+(2024a) alongside human annotators to assess data quality and filter inappropriate instances.
 4 Experiments
@@
 Baselines.
-To evaluate the effectiveness of our data construction, we select two existing multimodal safety alignment datasets for supervised fine-tuning as baselines that are listed as follows:.
-  • Direct: The original safety capability of the model without fine-tuning.
-  • VLGuard Zong et al. (2024): The first safety dataset specifically designed for fine-tuning of MLLMs.
-  • MIS Ding et al. (2025b): A multi-image safety dataset that combines visual perception with reasoning logic labels.
+To compare the effectiveness of our proposed dataset, we select three multimodal safety alignment datasets (MIS Ding et al.
+(2025b), VLGuard Zong et al.
+(2024), SPA-VL Zhang et al.
+(2024)) as our considered baselines.
+More details are listed in Appendix C.1.
 Training Details.
@@
 (2025) and LLaVA-CoT Xu et al.
-(2024) as base MLRMs for supervised fine-tuning, which clearly delineates the boundary between the thought process and the final response.
-We adopt supervised fine-tuning approach to the two models using the baseline datasets and our proposed dataset, respectively.
-The Details of the training procedure are provided in Appendix B.
+(2024) as base MLRMs for safety alignment training, which demonstrates competitive performance in multimodal reasoning tasks yet falls short in safety performance.
+We conduct comprehensive supervised fine-tuning on both models with all above datasets and our proposed one with detailed reasoning process.
+The details of the training configuration are provided in Appendix C.2.
 4.2 Results
-As shown in Table 5, the results demonstrate that the MLRMs achieve the best security performance after fine-tuning using our dataset on all benchmarks, compared with other baseline datasets.
-It is worth emphasizing that despite the complete absence of jailbreak attack data in our dataset, the proposed approach achieves a substantial reduction of ASR on jailbreak robustness benchmarks.
+As shown in Table 4, both R1-Onevision and LLaVA-CoT demonstrate improved safety alignment fine-tuned on TiS, substantially outperforming prior datasets.
+Specifically, our dataset TiS enhances safety performance on Figstep and SIUO by 30% compared to the best alternative baseline, effectively enabling the models to leverage their reasoning capabilities for deeper analysis and unsafe intention detection.
 Furthermore, since our dataset incorporates thought process that closely align with the data distribution used in MLRM training, the fine-tuned model preserved the original capability to generate coherent reasoning pathways.
 More case studies are provided in Appendix D.4.
-In addition, by fine-tuning the models respectively on the VLGuard, the MIS and the dataset we proposed, the safety performance of both models has been consistently improved, proving that well-designed constructing data is an effective solution to enhance the safety of the models.
+4.3 Discussion
+Ablation Studies.
+To investigate whether the proposed thought process can further enhance the safety performance of MLRMs, we conduct experiments by removing the thought component from TiS dataset, retaining only the answer portion.
+As shown in Table 4, retaining the thought process leads to more substantial improvements in safety performance compared to using answer-only data, with the exception of MM-SafetyBench.
+For certain instances of MM-SafetyBench, the model’s deeper reasoning leads it to categorize the input as neutral, which leads to the increase of ASR.The answer-only data also exhibits favorable performance, suggesting that the construction of more comprehensive and higher-quality datasets facilitates the safe fine-tuning of MLRMs.
+However, this can lead to generated responses that not only lack the thought process but also consist solely of brief replies such as “I’m sorry, I can’t assist it.”.More case studies are provided in Appendix D.3.
+Furthermore, this also demonstrates the feasibility of leveraging the reasoning capabilities of MLRMs to enhance the alignment with safety objectives.
+Analysis of Thought Process.
+To further assess the impact of Tis on enhancing the security of internal reasoning, we separately evaluate the safety of both thought process and final answer, as shown in Table 5.
+The results demonstrate the effectiveness of TiS in safeguarding both thought & answer response, with a substantial reduction in the ASR across all categories.
+Notably, TiS effectively reduces the proportion of cases in which both thought process and final answer are unsafe, decreasing from 42.80% to 0.20% on R1-Onevision, and from 54.60% to 0.40% on LLaVA-CoT.
+Furthermore, in scenarios where unsafe content appears solely in the intermediate thinking, our fine-tuning approach also achieves a marked improvement, reducing such instances from 14.20% to 2.60%.
+These results indicate that TiS largely enhances the reliability and security of the model’s internal reasoning process.
 5 Related Work
@@
 Safety of MLLMs.
-There are several studies that enhance the safety capabilities of MLLMs, which can be categorized into two types: the training-based method and the training-free method.
+Current studies enhancing the safety capabilities of MLLMs can be categorized into two types: the training-based method and the training-free method.
 The training-based method typically includes Supervised Fine-Tuning (SFT) Zong et al.
@@
 6 Conclusion
-In this work, we conduct a comprehensive evaluation and analysis of the safety of existing MLRMs, covering both jailbreak and comprehension benchmarks.
+In this work, we systematically evaluate and analyze the safety performance of existing MLRMs, covering both jailbreak robustness and safety-awareness benchmarks.
 The empirical results unveil several new findings, demonstrating that the safety performance of current MLRMs remains a significant concern.
@@
 Future research will focus on developing more efficient datasets and training strategies specifically designed for MLRMs.
+Limitations
+Our study primarily employs MLLMs as evaluative judges due to considerations of cost-efficiency and scalability.
+However, relying solely on MLLMs may compromise the accuracy of safety assessments, particularly in cases involving subtle forms of unsafe content or where the model fails to correctly interpret output response.
+Additionally, our evaluation includes only a collection of representative MLRMs, which does not capture the full diversity and reasoning capabilities of the broader range of available MLRMs.
 Appendix A Evaluation Details
+A.1 Benchmark and Evaluation Details
 Our evaluation encompasses five categories of safety benchmarks, which can be broadly divided into two types: jailbreak robustness and safety-awareness.
-Specifically, FigStep, MMSafetyBench, and JailBreaKV fall under the jailbreak robustness benchmark, for which we report the Attack Success Rate (ASR).
+Specifically, FigStep, MM-SafetyBench, and JailBreaKV fall under the jailbreak robustness benchmark, for which we report the Attack Success Rate (ASR).
 Meanwhile, MSSBench and SIUO are categorized as safety-awareness benchmarks, and we follow the respective safety metrics according to the original papers.
-The FigStep adopts the human evaluation as safety judgment.
-In our experiments, we follow the classification criteria described in their paper but apply an evaluation protocol consistent with that used for the MMSafetyBench.
-For the JailBreaKV, we follow the same evaluation settings as their work and utilize the Llama-Guard-Vision Chi et al.
+For the MM-SafetyBench, FigStep and SIUO, we employ GPT-4o-mini Hurst et al.
+(2024) and adopt identical configurations in teir work to ensure consistency across assessments.
+For the JailBreaKV, we follow the same evaluation settings of their work and utilize the Llama Guard 3 Vision Chi et al.
 (2024) as the evaluator.
+However, similar to the issues observed in MIS Ding et al.
+(2025b), we identify shortcomings in the safe category evaluation setup of MSSBench.
+As a result, we employed human evaluation as an alternative.
+Additional details are provided in Appendix A.2.
 The benchmarks are listed as follows:
@@
 $M$-SafetyBench.
-$M$-SafetyBench Liu et al.
+MM-SafetyBench Liu et al.
 (2024b) covers 13 categories and is primarily designed to evaluate the responses of MLLMs to malicious queries.
@@
 We report the evaluation metrics as defined in their paper.
-Appendix B Training Details
-For the training procedure, we utilize the LoRA Hu et al.
-(2022) to fine-tune the R1-Onevision and Llava-cot based on the LLaMA-Factory [^1] https://github.com/hiyouga/LLaMA-Factory framework.
-All training experiments are conducted using one NVIDIA A100-80G GPUs.
-The hyper-parameters for supervised fine-tuning are listed in Table 6.
-Hyper-Parameter Value
-lora rank 8
-learning rate 1.0e-5
-train epoch 1.5
-per_device_batchsize 1
-warm up ratio 0.1
-learning rate scheduler cosine
-Table 6: Hyper-parameters for supervised fine-tuning.
+Multimodal Large Reasoning Model Base Model
+LlamaV-o1 Thawakar et al. (2025) Llama 3.2-vision-11B
+LLaVA-CoT Xu et al. (2024) Llama 3.2-vision-11B
+Mulberry-Llama Yao et al. (2024) Llama 3.2-vision-11B
+LMM-R1 Peng et al. (2025b) Qwen2.5-VL-3B
+R1-Onevision Yang et al. (2025) Qwen2.5-VL-7B
+MM-Eureka Meng et al. (2025) InternVL-2.5-8B
+Kimi-VL-Thinking Team et al. (2025) Kimi-VL
+Skywork-R1V Peng et al. (2025a) -
+QVQ-Preview Team (2024) -
+Gemini2.0-Flash-Thinking DeepMind (2025) -
+Claude3.7-Sonnet-Thinking Anthropic (2025) -
+Table 6: List of the evaluated MLRMs and their corresponding base models.
+A.2 MSSBench Evaluation Details
+In the evaluation setup of MSSBench, LLM are required to classify responses into four safety levels.
+For certain examples in the safe category, responses that provide helpful advice but also emphasizing potential risks are evaluated as incorrect.
+As shown in Figure 6, the model provides the suggestion of using a heavy-duty drill while emphasizes safety as a priority.
+To address this limitation, we emplyed human evaluation as an alternative.
+The image, question, and response are delivered to three independent human annotators to obtain evaluation results.
+For the safe category, the evaluation primarily assesses whether the model exhibits unnecessary or excessive refusal.
+For the unsafe category, the evaluation focuses on whether the model can accurately detect and identify potential risks.
+Figure 6: A failure case of MSSBench evaluation with LLM-as-Judge method
+A.3 Model Details
+We evaluate a total of 11 MLRMs, Table 6 summarizes the Multimodal Large Reasoning Models evaluated for safety and their corresponding base models.
+Appendix B Dataset Details
+B.1 Detailed Pipeline
+Step 1: Data Preparation.
+Based on the categories provided in SPA-VL Zhang et al.
+(2024) and VLGuard Zong et al.
+(2024) datasets, we extract the raw data in the first step of construction for data preparation.
+In addition, we include the paired safety examples from VLGuard, where each image is matched with both a safe and an unsafe question.
+This design aims to improve the ability of the models to handle context-sensitive safety issues.
+However, the answers in SPA-VL are derived from pairwise preference data generated by large language models.
+Despite selecting the “chosen” responses as the final answers, a subset of them remain potential safety risks.
+The unsafe answers are subsequently screened and reconstructed.
+Step 2: Image Description Generation.
+To accommodate the input requirements of DeepSeek-R1 Guo et al.
+(2025), we convert images into caption form as input in advance.In particular, we employ Qwen2.5-VL-72B Bai et al.
+(2025) to generate image captions to ensure that the semantic loss during modality conversion is within an acceptable range in terms of security issues, with the prompt “Please provide a detailed description of this image.”.
+Step 3: Safety Thought Process Generation.
+After obtaining the image captions in the second step, the image caption, question, original response, and safety guidelines are provided to DeepSeek-R1 Guo et al.
+(2025) to generate safety-oriented thought process.
+The model first analyzes the question and the caption to obtain the underlying intent, then refers to the safety guidelines to generate a safety-oriented thought process corresponding to the provided response.
+Step 4: Thought Process Filter.
+The primary objective at this step is to filter and transform the original thought process.
+In particular, the caption field is processed to prevent potential misinterpretation by the MLRM.
+The original thought process from the previous step also contain redundant elements such as headers like “Safety Chain of Thought” which need to be removed and reformatted.
+Specific prompts used for this step are provided in the Figure 7.
 Model Chat Embodied AVG.
@@
 Llama 3.2-vision-11B(base) 85.62 22.07 53.85 90.79 10.53 50.66 52.26
-Llava-cot 93.67 12.67 53.17 97.37 1.32 49.34 51.26
+LLaVA-CoT 94.33 12.33 53.33 100.00 0.00 50.00 51.67
 Llama 3.2-vision-11B(base) 85.62 22.07 53.85 90.79 10.53 50.66 52.26
@@
 LMM-R1 83.00 33.00 58.00 69.70 26.32 48.03 53.02
-Qwen2.5-VL-7B(base) 93.33 11.00 52.17 93.42 6.58 50.00 51.09
-R1-Onevision 86.25 15.46 50.86 86.57 10.45 48.51 49.69
+Qwen2.5-VL-7B(base) 93.33 10.67 52.00 93.42 2.63 48.03 50.02
+R1-Onevision 86.60 15.12 50.86 86.57 7.46 47.01 48.94
 InternVL-2.5-8B (base) 92.00 11.33 51.67 98.68 1.32 50.00 50.84
 $M$-Eureka 90.67 11.67 51.17 97.37 2.63 50.00 50.59
-Kimi-VL-instruct (base) 91.67 8.67 50.17 100.00 1.43 50.71 50.44
+Kimi-VL (base) 91.67 8.67 50.17 100.00 1.43 50.71 50.44
 Kimi-VL-Thinking 83.33 23.67 53.5 98.67 0.00 49.33 51.42
 Table 7: Results on MSSBench in two distinct settings: chat and embodied scenarios.
-Appendix C Dataset Details
-C.1 Data Statistics
-Based on the categories provided in SPA-VL Zhang et al.
-(2024) and VLGuard Zong et al.
-(2024) datasets, we extract the raw data in the first step of construction for data preparation.
-In addition, we include the paired safety examples from VLGuard, where each image is matched with both a safe and an unsafe question.
-This design aims to improve the ability of the models to handle context-sensitive safety issues.
-However, the answers in SPA-VL are derived from pairwise preference data generated by large language models.
-Despite selecting the “chosen” responses as the final answers, a subset of them remain potential safety risks.
-The unsafe answers are subsequently screened and reconstructed.
+B.2 Dataset Statics
 The statistics of data instances in each category within our dataset after the final filtering process are shown in Table 8.
@@
 Economic harm 200
-Physical harm 200
+Physical harm 196
 Illegal activities 200
@@
 Safe 977
-All 4398
+All 4394
 Table 8: The statistics of our dataset
-C.2 Prompt for Data Conustruction
+B.3 Prompt for Data Construction
 This section presents the prompt and safety regulation used for safety thought process collection.
@@
 This strategy not only reduce the length of prompt but also enable tiered defensive measures for distinct safety-related issues.
+Appendix C Training Details
+C.1 Baselines
+To evaluate the effectiveness of our data construction, we select three existing multimodal safety alignment datasets as baselines that are listed as follows:.
+  • Direct: The original safety capability of the model without fine-tuning.
+  • VLGuard Zong et al. (2024): The first safety dataset specifically designed for fine-tuning of MLLMs.
+  • MIS Ding et al. (2025b): A multi-image safety dataset that combines visual perception with reasoning logic labels.
+  • SPA-VL Zhang et al. (2024): A safety preference alignment dataset for Vision Language Models.
+C.2 Traning Details
+For the training procedure, we utilize the LoRA Hu et al.
+(2022) to fine-tune the R1-Onevision and LLaVA-CoT based on the LLaMA-Factory [^1] https://github.com/hiyouga/LLaMA-Factory framework.
+For SPA-VL dataset, we use the DPO script for training.
+All training experiments are conducted using one NVIDIA A100-80G GPUs.
+The hyper-parameters for supervised fine-tuning are listed in Table 9.
+Hyper-Parameter Value
+lora rank 8
+learning rate 1.0e-5
+train epoch 1.5
+per_device_batchsize 1
+warm up ratio 0.1
+learning rate scheduler cosine
+Table 9: Hyper-parameters for supervised fine-tuning.
 Appendix D More Results
 D.1 Detailed Experiments on MSSBench
-In the evaluation protocol of MSS-Bench, scenarios are divided into chat and embodied settings, with safety scores calculated separately for safe and unsafe conditions.
+In the evaluation protocol of MSSBench, scenarios are divided into chat and embodied settings, with safety scores calculated separately for safe and unsafe conditions.
 The average safety scores are reported in Table 1, and the specific scores for each category are detailed in Table 7.
+D.2 Performance of Thought Process on MM-SafetyBench
+We also independently evaluate the safety performance of the thought process on MM-SafetyBench.
+This section presents the MLRMs’ performance and fine-tuned models’ performance in Table 10 and Table 11.
+  R1-Onevision Safe Answer Unsafe Answer
+Safe Thought 12.65% 1.90%
+Unsafe Thought 7.34% 78.11%
+Kimi-VL-Thinking Safe Answer Unsafe Answer
+Safe Thought 12.68% 2.74%
+Unsafe Thought 25.83% 58.75%
+LLaVA-CoT Safe Answer Unsafe Answer
+Safe Thought 15.89% 3.06%
+Unsafe Thought 11.85% 69.23%
+QVQ-Preview Safe Answer Unsafe Answer
+Safe Thought 24.52% 6.91%
+Unsafe Thought 6.19% 62.38%
+Skywork-R1V Safe Answer Unsafe Answer
+Safe Thought 19.88% 3.93%
+Unsafe Thought 7.44% 68.75%
+Table 10: Proportion of safety in the thought process and the final answer on MM-SafetyBench.
+  R1-Onevision Safe Answer Unsafe Answer
+Safe Thought 12.65% 1.90%
+Unsafe Thought 7.34% 78.11%
+R1-Onevision+Ours Safe Answer Unsafe Answer
+Safe Thought 74.58% 8.87%
+Unsafe Thought 16.55% 0.00%
+LLaVA-CoT Safe Answer Unsafe Answer
+Safe Thought 15.89% 3.06%
+Unsafe Thought 11.85% 69.23%
+LLaVA-CoT+Ours Safe Answer Unsafe Answer
+Safe Thought 71.73% 21.73%
+Unsafe Thought 6.43% 0.06%
+Table 11: Proportion of safety in the thought process and the final answer of fine-tuned model using TiS on MM-SafetyBench.
 Figure 7: Prompt used in data construction.
-D.2 Case Study
-Examples of responses generated by fine-tuned on VLGuard, MIS, and our dataset are illustrated in Figure 7.
+D.3 Case Study of Ablation Experiments
+Examples of responses generated by fine-tuned on TiS and TiS no thought dataset are illustrated in Figure 8 and Figure 9.
+The TiS dataset without thought is predominantly composed of unsafe examples.
+When fine-tuning MLRMs based on this data, the model tends to generate responses that consist solely of brief refusals.
+In particular, for safety-awareness examples, the model tends to either produce overly brief refusals or fail to detect unsafe risks.
+In contrary, the complete TiS data enables the detection of more unsafe intents during the thought process, while also providing more helpful information.
+Figure 8: Examples from R1-Onevision tuning on TiS w/o thought process. The red indicates the unsafe parts, while the blue indicates the safe parts.
+Figure 9: Examples from LLaVA-CoT tuning on TiS w/o thought process. The red indicates the unsafe parts, while the blue indicates the safe parts.
+D.4 Case Study
+Examples of responses generated by fine-tuned on VLGuard, MIS, SPA-VL and TiS dataset are illustrated in Figure 10.
 Our approach demonstrates the ability to retain the thought process of the models while decisively rejecting unsafe inputs and explicitly articulating the potential serious consequences associated with such queries.
-Figure 7: Examples of several baselines and our dataset on $M$-SafetyBench.
+Figure 10: Examples of several baselines and our dataset on MM-SafetyBench.
 We only highlight some fragments to show the representative difference between various methods.
 The red indicates the unsafe parts, while the blue indicates the safe parts.
+Figure 11: Examples of several baselines and our dataset on SIUO.
+We only highlight some fragments to show the representative difference between various methods.
+The red indicates the unsafe parts, while the blue indicates the safe parts.

### v4

@@
 The rapid development of Multimodal Large Reasoning Models (MLRMs) has demonstrated broad application potential, yet their safety and reliability remain critical concerns that require systematic exploration.
-To address this gap, we conduct a comprehensive and systematic safety evaluation of 11 MLRMs across 5 benchmarks and unveil prevalent safety degradation phenomena in most advanced models.
+To address this gap, we conduct a comprehensive and systematic safety evaluation of 13 MLRMs across 5 benchmarks and unveil prevalent safety degradation phenomena in most advanced models.
 Moreover, our analysis reveals distinct safety patterns across different benchmarks: significant safety degradation is observed across jailbreak robustness benchmarks, whereas safety-awareness benchmarks demonstrate less pronounced degradation.
@@
 To operationalize this insight, we construct a multimodal tuning dataset that incorporates a safety-oriented thought process.
-Experimental results from fine-tuning existing MLRMs with this dataset effectively enhances the safety on both jailbreak robustness and safety-awareness benchmarks.
+Experimental results from fine-tuning existing MLRMs with this dataset effectively enhance the safety on both jailbreak robustness and safety-awareness benchmarks.
 This study provides a new perspective for developing safe MLRMs.
-Our dataset is available at https://github.com/xinyuelou/Think-in-Safety.
+[^1] Our dataset is available at https://github.com/xinyuelou/Think-in-Safety.
 Warning: this paper contains example data that may be offensive or harmful.
 Think in Safety: Unveiling and Mitigating Safety Alignment Collapse in Multimodal Large Reasoning Model
-Xinyue Lou1111Equal contribution.
-Contact e-mail: louxinyue@bjtu.edu .cn, youlee@bjtu.edu.cn, You Li1111Equal contribution.
-Contact e-mail: louxinyue@bjtu.edu .cn, youlee@bjtu.edu.cn, Jinan Xu1, Xiangyu Shi1, Chi Chen2, Kaiyu Huang1 1 Beijing Jiaotong University, Beijing, China 2 Tsinghua University, Beijing, China
-22footnotetext: Corresponding author: kyhuang@bjtu.edu.cn
+Xinyue Lou1,2, You Li1,2, Jinan Xu1,2, Xiangyu Shi1,2, Chi Chen3, Kaiyu Huang1,2222Kaiyu Huang is the corresponding author.
+1Key Laboratory of Big Data & Artificial Intelligence in Transportation (Beijing Jiaotong University), Ministry of Education 2School of Computer Science and Technology, Beijing Jiaotong University 3Tsinghua University {louxinyue,kyhuang}@bjtu.edu.cn
 1 Introduction
@@
 Compared to unimodal LRMs, MLRMs exhibits the following research questions:
-RQ1: How does MLRMs affect safety on different types of benchmarks compared with Multimodal Large Language Models (MLLMs)?
+RQ1: How do MLRMs affect safety on different types of benchmarks compared with Multimodal Large Language Models (MLLMs)?
 RQ2: What are the risks of incorporating additional modalities in MLRMs for safety concerns?
@@
 […]
-Gemini2.0-Flash-Thinking 89.80 73.48 50.36 39.16 57.32
-Claude3.7-Sonnet-Thinking 32.20 33.94 0.49 59.24 57.76
-QVQ-Preview 70.80 69.29 37.86 34.13 50.68
-Skywork-R1V 85.80 72.68 35.28 35.33 50.42
-Llama 3.2-vision-11B(base) 55.80 38.45 5.71 37.13 52.26
-LlamaV-o1 59.40(+3.60) 53.93(+15.48) 13.57(+7.86) 33.93(-3.20) 51.59(-0.67)
-Llama 3.2-vision-11B(base) 55.80 38.45 5.71 37.13 52.26
-LLaVA-CoT 84.80(+29.00) 72.26(+33.81) 33.57(+27.86) 26.95(-10.18) 51.67(-0.59)
-Llama 3.2-vision-11B(base) 55.80 38.45 5.71 37.13 52.26
-Mulberry-Llama 67.40(+11.60) 64.70(+26.25) 13.21(+7.50) 37.72(+0.59) 54.09(+1.83)
-Qwen2.5-VL-3B(base) 66.27 66.18 12.14 24.55 52.35
-LMM-R1 69.80(+3.53) 68.15(+1.97) 18.21(+6.07) 21.56(-2.99) 53.02(+0.67)
-Qwen2.5-VL-7B(base) 67.20 66.49 12.50 29.94 50.02
-R1-Onevision 72.20(+5.00) 79.57(+13.08) 32.14(+19.64) 17.31(-12.63) 48.94(-1.08)
-InternVL-2.5-8B(base) 71.40 59.64 15.00 28.14 50.84
-$M$-Eureka 72.20(+0.80) 60.12(+0.48) 11.79(-3.21) 28.14(-0.00) 50.59(-0.25)
-Kimi-VL(base) 80.40 47.74 22.50 25.00 50.44
-Kimi-VL-Thinking 87.00(+6.60) 61.49(+13.75) 33.93(+11.43) 35.93(+10.93) 51.42(+0.98)
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 Table 1: Variation of safety performance for MLRMs across various benchmarks.
@@
 In contrast, jailbreak robustness benchmarks involve adversarial attacks, such as the inclusion of maliciously crafted prompt, aimed at circumventing the safety constraints of models.
-Given that these two task categories examine safety alignment from fundamentally different perspectives, we analyze these results separately.
+Given that these two task categories examine safety alignment from different perspectives, we analyze these results separately.
 Further details regarding the datasets are provided in Appendix A.1.
 Models and Configurations.
-We evaluate a total of 11 MLRMs, including both proprietary models and open-source models along with their corresponding base models, such as Kimi-VL-Thinking Team et al.
+We evaluate a total of 13 MLRMs, including both proprietary models and open-source models along with their corresponding base models, such as Kimi-VL-Thinking Team et al.
 (2025), R1-Onevision Yang et al.
@@
 Detailed information of the models are provided in Appendix A.3.
-For reasoning models that have undergone extensive fine-tuning—via either supervised learning or reinforcement learning—we additionally evaluate their corresponding base models.
+For reasoning models that have undergone extensive fine-tuning through either supervised learning or reinforcement learning, we additionally evaluate their corresponding base models.
 This approach allows us to trace safety-related changes across model variants.
-All experiments are conducted using four NVIDIA A100-80G GPUs.
+All experiments are conducted using two NVIDIA A100-80G GPUs.
 Metrics and Evaluator.
@@
 Further evaluation settings are detailed in Appendix A.1.
-Following the settings of MM-SafetyBench, MSS, and SIUO utilizing OpenAI API, we employ GPT-4o-mini Hurst et al.
+Following the settings of MM-SafetyBench Liu et al.
+(2024b), MSSBench Zhou et al.
+(2024), and SIUO Wang et al.
+(2025) utilizing OpenAI API, we employ GPT-4o-mini Hurst et al.
 (2024) as the evaluator and adopt identical configurations to ensure consistency across assessments.
@@
 2.2 Safety Results and Inspection
-Finding 1 ($\Rightarrow$ RQ1): The safety performance of MLRMs requires improvement and exhibits a notable security degradation phenomenon.
+Finding 1 ($\Rightarrow$ RQ1): The safety performance of MLRMs requires improvement and exhibits a notable safety degradation phenomenon.
 Overall Performance.
-As shown in Table 1, existing multimodal reasoning models exhibit significant security vulnerabilities, with most open-source models performing unsatisfactorily across various benchmarks.
+As shown in Table 1, existing multimodal reasoning models exhibit significant safety vulnerabilities, with most open-source models performing unsatisfactorily across various benchmarks.
 In particular, Kimi-VL-Thinking, Skywork-R1V and LLaVA-CoT successfully defend against only about 15% of malicious queries on FigStep.
 This finding highlights the necessity of ensuring safety alignment alongside improvements in multimodal reasoning capabilities.
-Security Degradation.
-Most reasoning models exhibit a notable increase in ASR on jailbreak robustness benchmarks relative to their corresponding foundation models, with an average increase of approximately 11.53%.
+Safety Degradation.
+Most reasoning models exhibit a notable increase in ASR on jailbreak robustness benchmarks relative to their corresponding foundation models, with an average increase of approximately 10.57%.
 This suggests that the internal safety alignment mechanisms are compromised during the process of enhancing reasoning capabilities, thereby rendering the models more susceptible to jailbreak attacks.
@@
 Finding 2 ($\Rightarrow$ RQ1): MLRMs paradoxically enhances safety performance via long reasoning on safety-awareness benchmarks, due to improved capability in identifying unsafe intent.
-Performance on Safety-Awareness.
-As shown in Table 1, the safety score of models such as Kimi-VL-Thinking Team et al.
-(2025), LMM-R1 Peng et al.
-(2025b), and MM-Eureka Meng et al.
-(2025) are higher than their corresponding base models on safety-awareness benchmark.
-The effect is particularly pronounced for Kimi-VL-Thinking model, which achieves a 10.9% improvement on the SIUO benchmark.
-As illustrated in Figure 2, there is a difference in focusing center according to the responses generated by these models and their corresponding base models.
-Reasoning models, when engaging in long reasoning, tend to focus more on potentially unsafe details present in both the visual and textual modalities.
-These models improve the safety performance through internal self-reflection and error-correction mechanisms.
-This encouraging finding offers a promising direction for mitigating safety degradation by leveraging the innate reasoning capabilities of the models themselves.
-Figure 3: Examples of self-deception in responses generated by MLRMs. The content of reframing intent of users is highlighted with yellow background, and harmful content is marked with the red font.
 […]
@@
 MMSafe is the abbreviation for $M$-SafetyBench.
+Figure 3: ASR scores of the thought process and the final answer generated by MLRMs on FigStep respectively.
+Performance on Safety-Awareness.
+As shown in Table 1, the safety score of models such as Kimi-VL-Thinking Team et al.
+(2025), Mixed-R1 Xu et al.
+(2025), and SophiaVL-R1  Fan et al.
+(2025) are higher than their corresponding base models on safety-awareness benchmark.
+The effect is particularly pronounced for SophiaVL-R1 model, which achieves an 11.38% improvement on the SIUO benchmark.
+As illustrated in Figure 2, there is a difference in focusing center according to the responses generated by these models and their corresponding base models.
+Reasoning models, when engaging in long reasoning, tend to focus more on potentially unsafe details present in both the visual and textual modalities.
+These models improve the safety performance through internal self-reflection and error-correction mechanisms.
+This encouraging finding offers a promising direction for mitigating safety degradation by leveraging the innate reasoning capabilities of the models themselves.
+Figure 4: Examples of self-deception in responses generated by MLRMs. The content of reframing intent of users is highlighted with yellow background, and harmful content is marked with the red font.
 Finding 3 ($\Rightarrow$ RQ2): The substitution of the image modality with text modality leads to a partial recovery of the safety defense capabilities of models.
@@
 In contrast, detecting harmful intent when reasoning over both image and text modalities presents greater challenges.
-Finding 4 ($\Rightarrow$ RQ3): MLRMs generate harmful content during the thought process, which poses a potential risk in response.
-Figure 4: ASR scores of the thought process and the final answer generated by MLRMs on FigStep respectively.
-Performance of Thought Process.
-To investigate the impact of reasoning pathways in MLRMs, we independently evaluate the safety performance of the thought process.
-As shown in Figure 4, the proportion of unsafe thought process generated by the four models (R1-Onevision, LLaVA-CoT, Kimi-VL-Thinking and Skywork-R1V) all exceeds 50 percent.
-Moreover, we analyze the thought process to identify how intermediate content contributes to the production of unsafe answers.
-As illustrated in Figure 3, although the model detects harmful intent in some cases, it does not reject the prompt outright.
-Instead, MLRM engages in self-deception by reframing the intent of users as educational, humorous, or otherwise benign, which rationalizes the model to generate harmful content.
-It is a critical challenge to address subtle intent rationalization in long reasoning, which needs to be explored in future work.
-Figure 5: Overview of our data construction pipeline.
-We propose a multi-stage pipeline to build the datasets based on various safety-related topics (Step 1: Word Cloud in Blue Block) and image description (Step 2: Red Block), which provides a thoughtful consideration of reasoning with safety, explicitly incorporating long CoT reasoning into the addressing process (Step 3: Green Block) and meticulously designed filtering mechanism (Step 4: Yellow Block).
-Finding 5 ($\Rightarrow$ RQ3): MLRMs attempt to eliminate unsafe content in the thought process via self-reflection, while there are also situations where unsafe answers are generated even if the thought process is secure.
   R1-Onevision Safe Answer Unsafe Answer
@@
 Table 3: Proportion of safety in the thought process and the final answer on FigStep.
+Finding 4 ($\Rightarrow$ RQ3): MLRMs generate harmful content during the thought process, which poses a potential risk in response.
+Performance of Thought Process.
+To investigate the impact of reasoning pathways in MLRMs, we independently evaluate the safety performance of the thought process.
+As shown in Figure 3, the proportion of unsafe thought process generated by the four models (R1-Onevision, LLaVA-CoT, Kimi-VL-Thinking and Skywork-R1V) all exceeds 50 percent.
+Moreover, we analyze the thought process to identify how intermediate content contributes to the production of unsafe answers.
+As illustrated in Figure 4, although the model detects harmful intent in some cases, it does not reject the prompt outright.
+Instead, MLRM engages in self-deception by reframing the intent of users as educational, humorous, or otherwise benign, which rationalizes the model to generate harmful content.
+It is a critical challenge to address subtle intent rationalization in long reasoning, which needs to be explored in future work.
+Figure 5: Examples of the safety inconsistency between the thought process and the final answer. The harmful content is marked with the red font.
+Figure 6: Overview of our data construction pipeline.
+We propose a multi-stage pipeline to build the datasets based on various safety-related topics (Step 1: Word Cloud in Blue Block) and image description (Step 2: Red Block), which provides a thoughtful consideration of reasoning with safety, explicitly incorporating long CoT reasoning into the addressing process (Step 3: Green Block) and meticulously designed filtering mechanism (Step 4: Yellow Block).
+Finding 5 ($\Rightarrow$ RQ3): MLRMs attempt to eliminate unsafe content in the thought process via self-reflection, while there are also situations where unsafe answers are generated even if the thought process is safe.
 Comparative Effect between Thought and Answer.
@@
 For example, the safe thought process leads to the unsafe answer in QVQ-Preview, with respect to 46.60% proportion.
-In specific examples, when the model meets with unsafe intents, it is not engaged in in-depth thinking to generate unsafe content in thought while directly generate unsafe answers in response.
-The inconsistency between safe thought and unsafe answer is due to the potential security risks inherent in the model.
+As shown in Figure 5(a), in specific examples, when the model encounters unsafe intents, the reasoning process focuses merely on the directivity for answering the question, without generating explicitly harmful actions.
+The inconsistency between safe thought and unsafe answer is due to the potential safety risks inherent in the model.
 On the contrary, the unsafe thought process also results in the safe final answer.
-It is attributed to the internal self-reflection and error-correction mechanisms of MLRMs, which enables models to eventually recognize the harmful nature of the intended prompt.
-The comparative results also provide potential insights for mitigating the issue of safety degradation in MLRMs.
+It is attributed to the internal self-reflection and error-correction mechanisms of MLRMs, which enables models to eventually recognize the harmful nature of the intended prompt as shown in Figure 5(b).
+The comparative results suggesting that we need to not only eliminate unsafe content in the reasoning process or answer section, but also ensure consistency between the thought process and the final answer.
 3 Data Construction
@@
 To address these issues, we propose a method for building a multimodal alignment dataset named TiS (Think in Safety), distinguished by its safety-oriented thought process and various safety-related topics.
-Through this method, we aim to leverage the innate reasoning capabilities of MLRM to improve its safety alignment.
+Through this method, we aim to leverage the innate reasoning capabilities of MLRMs to improve its safety alignment.
 As shown in Figure 6, we employ a multi-stage pipeline to construct our safety alignment dataset TiS.
@@
 After a filtering procedure, we finally obtain the dataset.
-To the best of our knowledge, TiS is the first safety dataset with the ability to retain reasoning chain for MLRMs.
+To the best of our knowledge, TiS is the first alignment dataset with the ability to retain reasoning chain for MLRMs.
 We provide a detailed illustration of the entire pipeline in Appendix B.1.
@@
 […]
-(R1-Onevision)
+[…]
 Direct 72.20 79.57 32.14 17.31 48.94
@@
 -w/o thought 20.80(-51.40) 16.37(-63.20) 0.00(-32.14) 63.47(+46.14) 63.92(+14.98)
-(LLaVA-CoT)
+[…]
 Direct 84.80 72.26 33.57 26.95 51.67
@@
 Due to the limitation of existing multimodal safety alignment datasets, we aim to construct a dataset enriched with safety-oriented thought process.
-However, considering the inherent security vulnerabilities of existing models and the resource constraints, constructing such a dataset entirely from scratch is largely impractical.
-Therefore, we augment existing datasets with short responses by adding structured and security-oriented thought process.
+However, considering the inherent safety vulnerabilities of existing models and the resource constraints, constructing such a dataset entirely from scratch is largely impractical.
+Therefore, we augment existing datasets with short responses by adding structured and safety-oriented thought process.
 Specifically, we select two multimodal safety alignment datasets as our original data sources: VLGuard Zong et al.
@@
 (2024).
-Rather than directly utilizing the entire datasets, we retain only the high-quality instances and ensure a balanced distribution across diverse safety-related topics.
+Rather than directly utilizing the entire datasets, we retain only a subset of instances to ensure a balanced distribution across diverse safety-related topics.
 Image Description Generation.
-To obtain a reliable thought process, we utilize the text-only reasoning model, Deepseek-R1 Guo et al.
-(2025), because instruction models often struggle to produce coherent reasoning, while MLRMs tend to suffer from alignment collapse under various conditions.
+To obtain a reliable thought process, we utilize the text-only reasoning model Deepseek-R1 Guo et al.
+(2025).
+This is because multimodal instruction models often struggle to produce coherent reasoning, while MLRMs tend to suffer from alignment collapse under various conditions.
 Meanwhile, close-source proprietary models are equipped with defensive safety guardrail, which hinders access to their internal reasoning processes particularly for safety-related questions.
@@
 (2024).
-Specific safety standards for various domains (law, medical, etc) are also clarified.
 Previous methods often require models to infer implicit safety rules from large volumes of training examples, which suffers from low efficiency and limited generalization.
-By contrast, our method provides sufficient safety standards for safety-centric reasoning content generation, facilitating deeper safety alignment and understanding rather than the first few tokens Qi et al.
-(2024).
+By contrast, we provide clear and structured safety guidelines to directly guide the model in generating safety-oriented thought process, encouraging MLRMs to think explicitly with safety considerations when producing responses.
+Besides, we develop the safety guidelines customized to the specific characteristics and risk profiles of each category.
 More details and prompt used in this step are provided in Appendix B.3.
 Thought Process Filtering.
-After obtaining the safety-oriented reasoning processes, we perform an additional modification and filtering step to ensure the quality and correctness of the data, making them better suited for safety enhancement.
+After obtaining the safety-oriented reasoning processes, we perform an additional modification and filtering step to ensure the quality and correctness of the data, making them better suited for safety alignment.
 This step addresses issues such as inaccuracies (e.g., responses that explicitly state “according to the caption”) and misalignments in reasoning.
@@
 4.2 Results
-As shown in Table 4, both R1-Onevision and LLaVA-CoT demonstrate improved safety alignment fine-tuned on TiS, substantially outperforming prior datasets.
-Specifically, our dataset TiS enhances safety performance on Figstep and SIUO by 30% compared to the best alternative baseline, effectively enabling the models to leverage their reasoning capabilities for deeper analysis and unsafe intention detection.
+As shown in Table 4, both R1-Onevision and LLaVA-CoT demonstrate improved safety alignment after fine-tuning on TiS, substantially outperforming prior datasets.
+Specifically, our dataset TiS enhances safety performance on FigStep and SIUO by at least 10% compared to the best alternative baseline, effectively enabling the models to leverage their reasoning capabilities for deeper analysis and unsafe intention detection.
 Furthermore, since our dataset incorporates thought process that closely align with the data distribution used in MLRM training, the fine-tuned model preserved the original capability to generate coherent reasoning pathways.
@@
 As shown in Table 4, retaining the thought process leads to more substantial improvements in safety performance compared to using answer-only data, with the exception of MM-SafetyBench.
-For certain instances of MM-SafetyBench, the model’s deeper reasoning leads it to categorize the input as neutral, which leads to the increase of ASR.The answer-only data also exhibits favorable performance, suggesting that the construction of more comprehensive and higher-quality datasets facilitates the safe fine-tuning of MLRMs.
+For certain instances of MM-SafetyBench, the model’s long reasoning leads it to categorize the input as neutral, which leads to the increase of ASR.The answer-only data also exhibits favorable performance, suggesting that the construction of more comprehensive and higher-quality datasets facilitates the safe fine-tuning of MLRMs.
 However, this can lead to generated responses that not only lack the thought process but also consist solely of brief replies such as “I’m sorry, I can’t assist it.”.More case studies are provided in Appendix D.3.
@@
 Analysis of Thought Process.
-To further assess the impact of Tis on enhancing the security of internal reasoning, we separately evaluate the safety of both thought process and final answer, as shown in Table 5.
-The results demonstrate the effectiveness of TiS in safeguarding both thought & answer response, with a substantial reduction in the ASR across all categories.
+To further assess the impact of Tis on enhancing the safety of internal reasoning, we separately evaluate the safety of both thought process and final answer, as shown in Table 5.
+The results demonstrate the effectiveness of TiS in safeguarding both thought and answer response, with a substantial reduction in the ASR across all categories.
 Notably, TiS effectively reduces the proportion of cases in which both thought process and final answer are unsafe, decreasing from 42.80% to 0.20% on R1-Onevision, and from 54.60% to 0.40% on LLaVA-CoT.
 Furthermore, in scenarios where unsafe content appears solely in the intermediate thinking, our fine-tuning approach also achieves a marked improvement, reducing such instances from 14.20% to 2.60%.
-These results indicate that TiS largely enhances the reliability and security of the model’s internal reasoning process.
+These results indicate that TiS largely enhances the reliability and safety of the model’s internal reasoning process.
 5 Related Work
 Safety of LRMs.
-With the rapid development and widespread deployment of LRMs, many works have paid attention to security of LRMs.
+With the rapid development and widespread deployment of LRMs, many works have paid attention to safety of LRMs.
 Several studies  Jiang et al.
-(2025); Parmar and Govindarajulu (2025) have conducted comprehensive security evaluations on LRMs, e.g., DeepSeek-R1 Guo et al.
+(2025); Parmar and Govindarajulu (2025) have conducted comprehensive safety evaluations on LRMs, e.g., DeepSeek-R1 Guo et al.
 (2025), revealing existing vulnerabilities of these models.
@@
 (2025) introduce a CoT jailbreak attack method, specifically targeting reasoning models.
-On the other hand, to mitigate the issues of security, recent studies Jiang et al.
+On the other hand, to mitigate the issues of safety, recent studies Jiang et al.
 (2025); Zhang et al.
@@
 Although SafeMLRM Fang et al.
-(2025) explores the security of MLRMs, the limitation of this work is that it only focuses on one security scenario, i.e., jailbreak robustness.
-The introduction of another modality in LRMs brings about significant shifts in model behavior, making it crucial to systematically assess the security challenges of MLRMs and develop effective defense methods.
+(2025) explores the safety of MLRMs and reveals three critical findings such as Reasoning Tax, the limitation of this work is that it only focuses on one safety scenario, i.e., jailbreak robustness.
+Our study is the first to investigate the safety behaviors of multimodal reasoning models in both jailbreak and awareness of safety scenarios and reveals more findings beyond Reasoning Tax.
 Safety of MLLMs.
@@
 For instance, MLLM-Protector Pi et al.
-(2024) offers a plug-and-play detector for harmful responses, and ETA Ding et al.
-(2025a) proposes an inference-time alignment framework to ensure safety compliance during deployment.
-In contrast to these methods, our study is the first to investigate the safety behaviors of reasoning models in both jailbreak and awareness of safety scenarios.
+(2024) offers a plug-and-play detector for harmful responses, ECSO Gou et al.
+(2024) converts the image inputs into text to exploit the safety of pre-aligned LLMs and ETA Ding et al.
+(2025a) proposes an inference-time alignment framework to ensure safety compliance.
+In contrast to these methods, our study is the first to propose potential solutions to address the issue of safety degradation of MLRMs.
 6 Conclusion
@@
 Additionally, our evaluation includes only a collection of representative MLRMs, which does not capture the full diversity and reasoning capabilities of the broader range of available MLRMs.
+Ethical Considerations
+In this paper, we primarily focus on investigation on the safety evaluation of MLRMs.
+All experiments are conducted using publicly released datasets in a controlled setting, thereby avoiding the creation or propagation of new harmful content.
+We highlight that the goal of our work is to reveal severe safety essue exhibited by MLRMs.
+Moreover, we designed TiS dataset to support the development of safer MLRMs without raising ethical concerns.
+Acknowledgments
+The research work descried in this paper has been supported by the National Nature Science Foundation of China (No.
+62376019, 62476023, 61976015, 61976016, 61876198 and 61370130), and the National Key R&D Program of China (2020AAA0108001).
+The authors would like to thank the anonymous reviewers for their valuable comments and suggestions to improve this paper.
 Appendix A Evaluation Details
@@
 For the MM-SafetyBench, FigStep and SIUO, we employ GPT-4o-mini Hurst et al.
-(2024) and adopt identical configurations in teir work to ensure consistency across assessments.
+(2024) and adopt identical configurations in their work to ensure consistency across assessments.
 For the JailBreaKV, we follow the same evaluation settings of their work and utilize the Llama Guard 3 Vision Chi et al.
@@
 (2024) is designed to evaluate the transferability of jailbreak techniques from LLMs to MLLMs, encompassing both text-based and image-based jailbreak attacks.
-Following the evaluation protocol of VLGuard, we conduct our experiments on the miniset provided in their paper, which contains 280 samples, following the setting of Hu et al.
-(2024).
+Following the setting of VLSBench Hu et al.
+(2024), we conduct our experiments on the miniset provided in their paper, which contains 280 samples.
 MSSBench.
@@
 R1-Onevision Yang et al. (2025) Qwen2.5-VL-7B
+Mixed-R1 Xu et al. (2025) Qwen2.5-VL-7B
+SophiaVL-R1 Fan et al. (2025) Qwen2.5-VL-7B
 MM-Eureka Meng et al. (2025) InternVL-2.5-8B
@@
 A.2 MSSBench Evaluation Details
-In the evaluation setup of MSSBench, LLM are required to classify responses into four safety levels.
+In the evaluation setup of MSSBench, LLMs are required to classify responses into four safety levels.
 For certain examples in the safe category, responses that provide helpful advice but also emphasizing potential risks are evaluated as incorrect.
 As shown in Figure 7, the model provides the suggestion of using a heavy-duty drill while emphasizes safety as a priority.
-To address this limitation, we emplyed human evaluation as an alternative.
+To address this limitation, we employed human evaluation as an alternative.
 The image, question, and response are delivered to three independent human annotators to obtain evaluation results.
@@
 A.3 Model Details
-We evaluate a total of 11 MLRMs, Table 6 summarizes the Multimodal Large Reasoning Models evaluated for safety and their corresponding base models.
+We evaluate a total of 13 MLRMs, Table 6 summarizes the Multimodal Large Reasoning Models evaluated for safety and their corresponding base models.
+Model Chat Embodied AVG.
+Safe Unsafe AVG. Safe Unsafe AVG.
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Table 7: Results on MSSBench in two distinct settings: chat and embodied scenarios.
 Appendix B Dataset Details
@@
 (2025), we convert images into caption form as input in advance.In particular, we employ Qwen2.5-VL-72B Bai et al.
-(2025) to generate image captions to ensure that the semantic loss during modality conversion is within an acceptable range in terms of security issues, with the prompt “Please provide a detailed description of this image.”.
+(2025) to generate image captions to ensure that the semantic loss during modality conversion is within an acceptable range in terms of safety issues, with the prompt “Please provide a detailed description of this image.”.
 Step 3: Safety Thought Process Generation.
@@
 Step 4: Thought Process Filter.
-The primary objective at this step is to filter and transform the original thought process.
-In particular, the caption field is processed to prevent potential misinterpretation by the MLRM.
-The original thought process from the previous step also contain redundant elements such as headers like “Safety Chain of Thought” which need to be removed and reformatted.
-Specific prompts used for this step are provided in the Figure 7.
-Model Chat Embodied AVG.
-Safe Unsafe AVG. Safe Unsafe AVG.
-Gemini2.0-Flash-Thinking 85.67 45.00 65.33 93.06 5.56 49.31 57.32
-Claude3.7-Sonnet-Thinking 75.84 45.97 60.91 93.42 15.79 54.61 57.76
-QVQ-Preview 80.67 24.67 52.67 53.95 43.42 48.68 50.68
-Skywork-R1V 86.67 15.00 50.83 98.68 1.32 50.00 50.42
-Llama 3.2-vision-11B(base) 85.62 22.07 53.85 90.79 10.53 50.66 52.26
-LlamaV-o1 59.33 47.00 53.17 100.00 0.00 50.00 51.59
-Llama 3.2-vision-11B(base) 85.62 22.07 53.85 90.79 10.53 50.66 52.26
-LLaVA-CoT 94.33 12.33 53.33 100.00 0.00 50.00 51.67
-Llama 3.2-vision-11B(base) 85.62 22.07 53.85 90.79 10.53 50.66 52.26
-Mulberry-Llama 81.33 35.00 58.17 97.37 2.63 50.00 54.09
-Qwen2.5-VL-3B(base) 83.67 31.00 57.33 59.21 35.53 47.37 52.35
-LMM-R1 83.00 33.00 58.00 69.70 26.32 48.03 53.02
-Qwen2.5-VL-7B(base) 93.33 10.67 52.00 93.42 2.63 48.03 50.02
-R1-Onevision 86.60 15.12 50.86 86.57 7.46 47.01 48.94
-InternVL-2.5-8B (base) 92.00 11.33 51.67 98.68 1.32 50.00 50.84
-$M$-Eureka 90.67 11.67 51.17 97.37 2.63 50.00 50.59
-Kimi-VL (base) 91.67 8.67 50.17 100.00 1.43 50.71 50.44
-Kimi-VL-Thinking 83.33 23.67 53.5 98.67 0.00 49.33 51.42
-Table 7: Results on MSSBench in two distinct settings: chat and embodied scenarios.
+The primary objective in this step is to filter and transform the original thought process.
+In particular, the caption field is processed to prevent potential misinterpretation by the MLRMs.
+The original thought process from the previous step also contains redundant elements such as headers like “Safety Chain of Thought” which need to be removed and reformatted.
+Specific prompts used for this step are provided in Figure 8.
 B.2 Dataset Statics
@@
 During the collection of safety thought process data, we supplied the model with category-specific safety regulations derived from the raw data classification.
-This strategy not only reduce the length of prompt but also enable tiered defensive measures for distinct safety-related issues.
+This strategy not only reduces the length of prompt but also enables tiered defensive measures for distinct safety-related issues.
 Appendix C Training Details
@@
 (2022) to fine-tune the R1-Onevision and LLaVA-CoT based on the LLaMA-Factory [^1] https://github.com/hiyouga/LLaMA-Factory framework.
-For SPA-VL dataset, we use the DPO script for training.
-All training experiments are conducted using one NVIDIA A100-80G GPUs.
+For the SPA-VL dataset, to ensure fairness, we randomly selected 4,000 data samples with balanced categories and used the DPO script for training.
+All training experiments are conducted using one NVIDIA A100-80G GPU.
 The hyper-parameters for supervised fine-tuning are listed in Table 9.
@@
   R1-Onevision Safe Answer Unsafe Answer
-Safe Thought 12.65% 1.90%
-Unsafe Thought 7.34% 78.11%
+Safe Thought 13.71% 41.66%
+Unsafe Thought 6.72% 37.91%
 Kimi-VL-Thinking Safe Answer Unsafe Answer
-Safe Thought 12.68% 2.74%
-Unsafe Thought 25.83% 58.75%
+Safe Thought 13.27% 40.65%
+Unsafe Thought 25.25% 20.83%
 LLaVA-CoT Safe Answer Unsafe Answer
-Safe Thought 15.89% 3.06%
-Unsafe Thought 11.85% 69.23%
+Safe Thought 18.69% 40.77%
+Unsafe Thought 9.05% 31.49%
 QVQ-Preview Safe Answer Unsafe Answer
-Safe Thought 24.52% 6.91%
-Unsafe Thought 6.19% 62.38%
+Safe Thought 23.15% 44.76%
+Unsafe Thought 7.50% 24.52%
 Skywork-R1V Safe Answer Unsafe Answer
-Safe Thought 19.88% 3.93%
-Unsafe Thought 7.44% 68.75%
+Safe Thought 21.07% 42.26%
+Unsafe Thought 6.25% 30.42%
 Table 10: Proportion of safety in the thought process and the final answer on MM-SafetyBench.
   R1-Onevision Safe Answer Unsafe Answer
-Safe Thought 12.65% 1.90%
-Unsafe Thought 7.34% 78.11%
-R1-Onevision+Ours Safe Answer Unsafe Answer
-Safe Thought 74.58% 8.87%
-Unsafe Thought 16.55% 0.00%
+Safe Thought 13.71% 41.66%
+Unsafe Thought 6.72% 37.91%
+R1-Onevision+TiS Safe Answer Unsafe Answer
+Safe Thought 77.32% 21.49%
+Unsafe Thought 0.95% 0.24%
 LLaVA-CoT Safe Answer Unsafe Answer
-Safe Thought 15.89% 3.06%
-Unsafe Thought 11.85% 69.23%
-LLaVA-CoT+Ours Safe Answer Unsafe Answer
-Safe Thought 71.73% 21.73%
-Unsafe Thought 6.43% 0.06%
+Safe Thought 18.69% 40.77%
+Unsafe Thought 9.05% 31.49%
+LLaVA-CoT+TiS Safe Answer Unsafe Answer
+Safe Thought 84.82% 8.75%
+Unsafe Thought 5.89% 0.54%
 Table 11: Proportion of safety in the thought process and the final answer of fine-tuned model using TiS on MM-SafetyBench.
-Figure 7: Prompt used in data construction.
+Figure 8: Prompts used in data construction.
 D.3 Case Study of Ablation Experiments
-Examples of responses generated by fine-tuned on TiS and TiS no thought dataset are illustrated in Figure 8 and Figure 9.
+Examples of responses generated by models fine-tuned on TiS and TiS no thought dataset are illustrated in Figure 9 and Figure 10.
 The TiS dataset without thought is predominantly composed of unsafe examples.
@@
 In particular, for safety-awareness examples, the model tends to either produce overly brief refusals or fail to detect unsafe risks.
-In contrary, the complete TiS data enables the detection of more unsafe intents during the thought process, while also providing more helpful information.
+In contrast, the complete TiS data enables the detection of more unsafe intents during the thought process, while also providing more helpful information.
 Figure 9: Examples from R1-Onevision tuning on TiS w/o thought process. The red indicates the unsafe parts, while the blue indicates the safe parts.
@@
 D.4 Case Study
-Examples of responses generated by fine-tuned on VLGuard, MIS, SPA-VL and TiS dataset are illustrated in Figure 10.
+Examples of responses generated models fine-tuned on VLGuard, MIS, SPA-VL and TiS dataset are illustrated in Figure 11 and Figure 12.
 Our approach demonstrates the ability to retain the thought process of the models while decisively rejecting unsafe inputs and explicitly articulating the potential serious consequences associated with such queries.